### v1

# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

Abstract
Generative AI systems like foundation models (FMs) must align well with human values to ensure their behavior is helpful and trustworthy. While Reinforcement Learning from Human Feedback (RLHF) has shown promise for optimizing model performance using human judgments, existing RLHF pipelines predominantly rely on immediate feedback, which can fail to accurately reflect the downstream impact of an interaction on users’ utility. We demonstrate that feedback based on evaluators’ foresight estimates of downstream consequences systematically induces Goodhart’s Law dynamics, incentivizing misaligned behaviors like sycophancy and deception and ultimately degrading user outcomes. To alleviate this, we propose decoupling evaluation from prediction by refocusing RLHF on hindsight feedback. Our theoretical analysis reveals that conditioning evaluator feedback on downstream observations mitigates misalignment and improves expected human utility, even when these observations are simulated by the AI system itself. To leverage this insight in a practical alignment algorithm, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which first simulates plausible consequences and then elicits feedback to assess what behaviors were genuinely beneficial in hindsight. We apply RLHS to two widely-employed online and offline preference optimization methods—Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO)—and show empirically that misalignment is significantly reduced with both methods. Through an online human user study, we show that RLHS consistently outperforms RLHF in helping users achieve their goals and earns higher satisfaction ratings, despite being trained solely with simulated hindsight feedback. These results underscore the importance of focusing on long-term consequences, even simulated ones, to mitigate misalignment in RLHF.
1 Introduction
Aligning artificial intelligence (AI) systems with human values and intentions is crucial to ensuring they behave in ways that are helpful, honest, and trustworthy. A widely-deployed method for achieving this alignment is through human feedback (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it often relies heavily on human perceptions during interactions, which may not accurately reflect the downstream consequences of the service provided. Such myopic feedback can misguide the model’s behavior and limit its effectiveness in aligning with human values. For example, human evaluators could misjudge an interaction on the spot, due to limited resources (e.g., partial observability; Casper et al. 2023; Lang et al. 2024) or limited bandwidth (e.g., constraints on time, attention, or care; Pandey et al. 2022; Chmielewski & Kucker 2020), leading to incomplete or misinformed feedback. A recent study has theorized that partial observability of an AI assistant’s task execution during human–AI interaction can lead RLHF to learn deceptive behaviors (Lang et al., 2024).
Figure 1: RLHF can incentivize AI systems to provide inaccurate or deceptive information to their users, prioritizing positive on-the-spot feedback and neglecting long-term consequences. For example, a customer may prefer to hear good news while shopping but will ultimately be disappointed (and objectively worse off) if stuck with an ill-informed purchase. The proposed RLHS introduces hindsight for human feedback, focusing on evaluations after simulating the outcome. This enables more informed feedback that improves alignment between the AI and the human’s true utility.
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. We hypothesize that relying on human users’ prediction of the helpfulness of an interaction right after it takes place creates a pernicious Goodhart’s law dynamic: incentivizing the AI system to increase users’ subjective foresight value will favor inducing unrealistically optimistic expectations in users—while at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded outcomes.
We provide substantial empirical evidence that indeed this phenomenon can arise even in simple settings: we find that immediate human feedback elicited at the end of the human–AI interaction frequently misrepresents true utility in consultancy-type interactions, and, when used as a proxy for it in RLHF fine-tuning, it systematically drives misalignment with human goals (Fig. 1, top). Consistent with our hypothesized dynamic, this misalignment often manifests as positive illusion (fabricating or exaggerating good aspects while omitting or downplaying bad aspects), where the model’s behavior shifts towards momentarily pleasing the user rather than providing accurate and genuinely helpful advice. This consistently leads users to make ill-informed decisions whose poor downstream outcomes contrast starkly with their high satisfaction rating at the end of the interaction.
To address these open challenges, we propose to leverage hindsight as a simple but effective misalignment mitigation mechanism, in which evaluators experience the downstream outcomes of an interaction before being asked for feedback on the model. We provide both theoretical analysis and extensive empirical studies to show the efficacy of hindsight in significantly reducing misalignment of RLHF-trained models. To circumvent the material and ethical difficulties in exposing real people to real consequences, we introduce a novel alignment algorithm called Reinforcement Learning from Hindsight Simulation (RLHS), an alternative to RLHF that uses an AI world model during training to simulate likely human decisions and their downstream outcomes. Despite lacking any privileged information on real outcomes, the method effectively relieves evaluators of the need to predict the future consequences of an AI output—thus removing the AI incentive to influence these predictions.
Our key finding is that equipping evaluator feedback with the benefit of hindsight—even if this is simulated using imperfect models—can significantly reduce model misalignment with the evaluator’s true utility, decreasing the chances of deceptive and misleading outputs. We implement hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and show empirically that it greatly improves alignment in both training paradigms. We also present results from human user studies, in which RLHS consistently improves both users’ ground-truth utility and subjective satisfaction, despite being trained with only simulated hindsight feedback. Our comparative findings demonstrate that RLHS significantly outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for real human feedback, and has been shown to produce results closely resembling that of RLHF (Bai et al., 2022b; Lee et al., 2023).
2 Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We can model such a problem as a partially observable Markov decision process (POMDP) defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$ encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
AI-Assisted Human Decision-Making. When the human consults an AI system (e.g., a FM) to help with their decision problem, we may augment the above problem with the human–AI interaction. The resulting Assisted POMDP is a tuple $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}=(\mathcal{S},\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\mathcal{O}^{\textit{H}},\mathcal{O}^{\textit{AI}},\mathcal{T},{O}^{\textit{H}},{O}^{\textit{AI}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ and $\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ are the sets of interactive actions available to the human and AI system, $\mathcal{O}^{\textit{AI}}$ is the AI’s observation space, and ${O}^{\textit{AI}}$ is the AI’s observation map ${O}^{\textit{AI}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{AI}})$. In this model, the AI takes an advisory role: it can respond to a human’s interactive action ${a}_{\rightleftharpoons}^{\textit{H}}\in\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ (e.g., a query through a chat interface) with its own ${a}^{\textit{AI}}_{\rightleftharpoons}\in\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ (e.g., a generated text or multimedia output). After one or multiple rounds of such interactions, the human may take a physical action ${a}^{\textit{H}}\in\mathcal{A}^{\textit{H}}$ to affect the evolution of the world state ${s}$. This Assisted POMDP is a special case of a partially observable stochastic game (POSG) (Hansen et al., 2004). In such interactions, the AI’s goal is to influence the human’s internal state ${z}^{\textit{H}}$ towards maximizing the rewards $r({s},{a}^{\textit{H}};\theta)$ accrued over time. This, however, is made challenging by the AI’s fundamental uncertainty about the human’s preferences ${\theta}^{\textit{H}}$.
Reinforcement Learning from Human Feedback (RLHF). RLHF aims to learn the human’s preferences ${\theta}^{\textit{H}}$ from human feedback data, which typically involves three key steps. In Step 1, the human is asked to provide feedback on some state sequences ${\mathbf{{s}}}=({s}_{0},{s}_{1},\ldots,{s}_{T})$ (e.g., a human–AI dialogue), with ${s}_{t}\in\mathcal{S},~{}\forall t=0,1,\ldots,T$. For example, in binary comparison (Christiano et al., 2017), assuming human is a Boltzmann rational decision maker (Luce, 1959), the probability that the human prefers ${\mathbf{{s}}}$ over ${\mathbf{{s}}}^{\prime}$ is $P^{r}_{T}({\mathbf{{s}}}\succ{\mathbf{{s}}}^{\prime})=\sigma(\beta(R_{T}({\mathbf{{s}}})-R_{T}({\mathbf{{s}}}^{\prime})))$, where $\sigma(\cdot)$ is the sigmoid function, $\beta>0$ is the inverse temperature parameter, and $R_{T}({\mathbf{{s}}})=\sum_{t=0}^{T}\gamma^{t}r({s}_{t})$ is the return received by state sequence ${\mathbf{{s}}}$.. Step 2 is to fit a reward function $\hat{r}$ based on a dataset containing state sequences paired with human feedback, aiming for $\hat{r}$ to resemble $r$ as closely as possible. Step 3 is to compute an AI policy $\hat{{\pi}}:\mathcal{S}\rightarrow\Delta(\mathcal{A}^{\textit{AI}}_{\rightleftharpoons})$ that maximizes the return based on the estimated reward $\hat{r}$, i.e., $\hat{{\pi}}=\operatorname*{arg\,max}_{\pi}{U}_{T}({\pi})$, where ${U}_{T}({\pi}):=\operatorname*{\mathbb{E}}_{{\mathbf{{s}}}\sim p^{\pi}}[\hat{R}_{T}({\mathbf{{s}}})]$ is the expected utility of ${\pi}$, and $p^{\pi}$ is the on-policy distribution of state sequence ${\mathbf{{s}}}$ under $P_{0}$, $\mathcal{T}$, and ${\pi}$. Due to the lack of an analytical model for $\mathcal{T}$ and the high-dimensional nature of aligning modern AI models, reinforcement learning (RL) is often used to approximately optimize the policy at scale. Recent studies have revealed that RLHF can lead to misalignment when the human gives feedback based on partial observations ${\mathbf{{o}}}^{\textit{H}}=({o}^{\textit{H}}_{0},{o}^{\textit{H}}_{1},\ldots,{o}^{\textit{H}}_{T})$ rather than the previously assumed—but rarely realistic—full state sequences, resulting in deceptive or manipulative AI behaviors (Casper et al., 2023; Lang et al., 2024). We argue that RLHF misalignment more generally emerges in settings with significant human uncertainty, whether perceptual, predictive, or a combination of the two. We propose to take advantage of the general insight that assessments about past outcomes that the evaluator has experienced would be significantly less uncertain (and thus less influenceable) than assessments about future outcomes that are yet to unfold.
3 Alignment Algorithm: RL from Hindsight Simulation
To address the misalignment caused by human uncertainty in RLHF, we introduce Reinforcement Learning from Hindsight Simulation (RLHS). Our central contention is that by decoupling human feedback on the downstream outcomes of an interaction from the prediction of these outcomes, the learned human reward model and corresponding AI policy will be substantially better aligned.
3.1 Hindsight Mitigates Misalignment
Given a predictive model of the human, the AI’s decision problem in the Assisted POMDP game $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}$ in Section 2 can be reformulated as a POMDP $\mathcal{P}^{\textit{AI}}_{\rightleftharpoons}=(\bar{\mathcal{S}},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\bar{\mathcal{O}}^{\textit{AI}},\bar{\mathcal{T}},\bar{O}^{\textit{AI}},\bar{P}_{0},\bar{r},\gamma)$, where $\bar{\mathcal{S}}=\mathcal{S}\times\Theta^{\textit{H}}\times{\mathcal{Z}^{\textit{H}}}$, $\bar{\mathcal{O}}^{\textit{AI}}=\mathcal{O}^{\textit{AI}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$, $\bar{\mathcal{T}}=(\mathcal{T},\mathcal{T}_{\theta},\mathcal{T}^{\textit{H}})$, $\bar{P}_{0}\in\Delta(\bar{\mathcal{S}})$, and $\bar{r}({s},{z}^{\textit{H}},{\theta}^{\textit{H}})=\operatorname*{\mathbb{E}}_{{a}^{H}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}})}r({s},{a}^{\textit{H}};{\theta}^{\textit{H}})$. Here, $\mathcal{T}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\times\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}\rightarrow{\mathcal{Z}^{\textit{H}}}$ is the transition kernel of the human’s internal state, modeling how the human’s knowledge about the world state is evolved based on new observations and interactions with the AI; we treat ${\theta}^{\textit{H}}$ as a constant for the purposes of this paper, with $\mathcal{T}_{\theta}$ as the identity map. Finally, $\pi^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\to\Delta(\bar{\mathcal{A}}^{\textit{H}})$, with $\bar{\mathcal{A}}^{\textit{H}}:=\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$. In practice the human model can be a black box (e.g., a web-trained FM). Due to the complexity of POMDP $\mathcal{P}^{\textit{AI}}_{\rightleftharpoons}$, we aim to solve it approximately using RL with hindsight feedback provided by an evaluator, which we explain in detail below.
Since the human’s utility is inherited from their original decision problem $\mathcal{P}^{\textit{H}}$, the expected utility generated by an AI policy ${\pi}^{\textit{AI}}$ is the expected return achieved by the human’s course of action. For the purposes of RLHF, we can assume that the human begins taking physical actions after the interaction:
${U}^{\textit{H}}({\pi}^{\textit{AI}}):=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\bar{s}_{0}\sim\bar{P}_{0},\quad\,\mathcal{T}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{\tau},~{}{a}^{\textit{AI}}_{{\rightleftharpoons},\tau}),~{}\\ {a}^{\textit{H}}_{\tau}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{\tau}),\;{a}^{\textit{AI}}_{{\rightleftharpoons},\tau}\sim{\pi}^{\textit{AI}}(\cdot\mid z^{\textit{AI}}_{\tau})\end{subarray}}\left[\sum_{t=T+1}^{\infty}\gamma^{t}r({s}_{t},{a}^{H}_{t};{\theta}^{\textit{H}})\right],$ (1)
where $t=0,1,\ldots,T$ is the human–AI interaction phase and $t=T+1,T+2,\ldots,$is the downstream human acting phase. If the human evaluates their utility immediately after the human–AI interaction phase, relying only on predictions, this is the foresight value. If, instead, they wait until after the human acting phase and use realized outcome, this is the hindsight value. The following sections define these concepts formally.
Definition 1 (Hindsight Value).
The hindsight value assessed by the human at time $k\geq 0$ is equal to the expected return received so far given the human’s available information at time $k$. In this paper we will assume that the human can accurately estimate all rewards received so far, i.e.,
$\overleftarrow{{V}}_{k}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{\begin{subarray}{c}{s}_{t+1}\sim\mathcal{T}(\cdot|{s}_{t},{a}^{\textit{H}}_{t}),~{}{a}^{\textit{H}}_{t}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t})\end{subarray}}\!\biggl{[}\sum_{t=0}^{k}\gamma^{t}\,r({s}_{t},{a}^{\textit{H}}_{t})\biggr{]},$ (2)
where $\mathcal{T}$ is the environment transition kernel, which is unaffected by the AI’s output.
Definition 2 (Foresight Value).
The foresight value at time $k\geq 0$ is the expected reward-to-go from $k$ to $k+N$ under the human’s subjective prediction of future outcomes. Foresight uses the human’s own internal predictive model of what states and actions might arise. Formally:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{\begin{subarray}{c}{s}_{t}\sim P(\cdot\mid{z}^{\textit{H}}_{t}),~{}{a}^{\textit{H}}_{t}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t})\end{subarray}}\left[\sum_{t=k}^{k+N}\gamma^{t}r({s}_{t},{a}^{\textit{H}}_{t})\right],$ (3)
where $P(\cdot\mid{z}^{\textit{H}}_{t})$ is the human’s internal predictive distribution over future states. Neither the states ${s}_{t}$ nor the rewards here are necessarily generated by the environment’s true transition kernel; rather, they are drawn from the human’s imaginative predictions.
Figure 2: Illustration of hindsight’s advantage: Delaying human feedback until the human has experienced the outcome corresponding to the bulk of reward significantly mitigates the misalignment in the AI’s learned reward model.
Using this foresight-based feedback, the human evaluator relies on predictions of future actions and states, which can be influenced by the AI. Consequently, the predicted human utility using foresight is: $\tilde{{U}}^{\textit{H}}_{T+N}({\pi}^{\textit{AI}})\;=\;\overrightarrow{{V}}_{T+N}({z}^{\textit{H}}_{T}).$ However, RL from Hindsight Simulation (RLHS) aims to reduce the human’s reliance on such predictions by shifting the evaluation from foresight to hindsight value. By simulating the downstream acting phase for $N$ steps, the human can directly observe realized outcomes rather than guess them, thus: ${U}^{\textit{H}}_{T+N}({\pi}^{\textit{AI}})\;=\;\overleftarrow{{V}}_{T+N}({z}^{\textit{H}}_{T}).$
In hindsight-based evaluation, the states $\{{s}_{t}\}$ come from actual world rollouts rather than predictions conditioned on the human’s internal state which can be influenced by the AI. This makes hindsight feedback substantially more robust and less prone to reward hacking, as it depends on concrete observed outcomes rather than speculative future scenarios influenced by the AI’s advice.
In the following, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility. Consider an oracle aligned AI policy ${\pi}^{*}$ that operates knowing the human preference ${\theta}^{\textit{H}}$. The following lemma establishes that, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$, the difference in finite-hindsight utility estimation becomes an exponentially accurate estimate of the difference in true utility as the hindsight horizon $N$ increases.
Lemma 1.
Let the finite hindsight utility estimate ${U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})$ be the $N$-step truncation of the expected utility sum in equation 1, and let the reward function $r$ be bounded by $\underline{r}\leq r({s},{a}^{\textit{H}})\leq\bar{r}$ for all ${s}\in\mathcal{S}$, ${a}^{\textit{H}}\in\mathcal{A}^{H}$, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$. Then, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$,
${U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})-{U}_{N}^{\textit{H}}(\tilde{\pi}^{\textit{AI}})\in\mathcal{B}\Big{(}{U}^{\textit{H}}({\pi}^{\textit{AI}})-{U}^{\textit{H}}(\tilde{\pi}^{\textit{AI}}),\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\,.$
Proof.
The lemma follows directly from bounding the tail of the series from term $T+N+1$. ∎
Applying the same logic of this lemma to individual executions and assuming a Boltzmann-rational evaluator like the one discussed in Section 2 (and often considered for theoretical purposes when analyzing RLHF methods), we obtain the following result.
Theorem 1.
Suppose the human evaluator is presented a finite-horizon hindsight of $N$ steps and makes Boltzmann-rational binary preference choices with inverse temperature $\beta$. Then the probability that the human prefers a hindsight observation ${\mathbf{{o}}}_{0:T+N}$ over another $\bar{\mathbf{{o}}}_{0:T+N}$ from the same initial information state $P({\mathbf{{o}}}_{0:T+N}\succ\bar{\mathbf{{o}}}_{0:T+N})$ is within the range
$\sigma\left(\beta\Big{(}R_{T}({\mathbf{{o}}}_{0:T+N})-R_{T}(\bar{\mathbf{{o}}}_{0:T+N})\pm\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\right).$
This ensures that, for a sufficiently large hindsight horizon, the hindsight feedback of a Boltzmann-rational human evaluator can be made arbitrarily close—in probability—to the ideal infinite-horizon oracle feedback. We view this as providing theoretical support for the empirically observed value of hindsight with respect to default RLHF (which corresponds to the degenerate case $N=0$).
3.2 Implementation: Hindsight Simulation with AI Feedback
Figure 3: Qualitative results for Llama-2-7b trained with either immediate feedback (RLHF) or partial hindsight (RLHS). The RLHF model (trained with immediate feedback) deceives the user by falsely claiming Options A and C meet the customer’s 8K resolution requirement, though neither does. In contrast, the RLHS model truthfully states that none of the options include 8K resolution.
Hindsight Simulation. While we have demonstrated theoretically that providing hindsight can mitigate misalignment in RLHF, exposing humans to real consequences can circumvent material and ethical difficulties. To address this, we introduce the concept of hindsight simulation—the namesake of our core contribution, RLHS —which allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve collecting feedback from human evaluators or employing another language model as a proxy to simulate the feedback process. After an evaluator makes a decision based on their interaction with the AI (e.g., purchasing an item), the system provides ground truth information about the outcome, i.e., the hindsight (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the decision’s outcome and their prior interaction with the model.
This feedback typically takes the form of a rating or binary preference, similar to the feedback used in conventional RLHF. However, unlike the immediate feedback provided solely during an interaction without access to the decision’s consequences, feedback obtained through hindsight simulation is more informed as it incorporates long-term outcomes. This aligns with the reasoning presented in Section 3.1 and demonstrates the potential for improving alignment through simulated hindsight.
We implement this approach with two subroutines: (i) partial hindsight, where only a limited set of hindsight information is available to the agent, in a way that more closely matches real-world scenarios, and (ii) oracle hindsight, where the agent has access to full set of hindsight information. In our empirical studies employing both partial and oracle hindsight, we provide insights into how extending the hindsight step (i.e., revealing additional outcome information to the agent) can improve the alignment performance of the model.
Illustrative Example: Marketplace Chatbot. We demonstrate the practical impact of RLHS by applying it to a marketplace AI chatbot. The chatbot’s goal is to assist customers in making purchasing decisions by providing recommendations based on available product information. We assume that both customers and the chatbot have access to some public information, such as a list of items and their prices, but customers have their internal preferences, e.g., wanting a TV with 8K resolution, that are unknown to the chatbot. To the best of our knowledge, existing RLHF schemes deployed for training marketplace chatbots (e.g., Amazon, 2024) use customer feedback solely based on the interaction (i.e., if they feel happy about the chatbot’s service) but not on the outcome (i.e., if the purchased item actually meets their preferences), potentially causing misalignment.
Our proposed hindsight simulation approach aims to mitigate this issue by deferring the humans’ feedback until they have been informed of the outcome of their decisions, e.g., they have received the product and verified that their expectations are (not) met. In hindsight simulation, the simulated customer interacts with the chatbot, makes a purchasing decision, checks the outcome (hindsight) provided by the system, and provides feedback on the customer’s satisfaction with the service.
4 Experimental Design
4.1 Data Collection
Preference Data Collection. Our training data collection process closely follows the standard RLHF data collection pipeline (Stiennon et al., 2020; Ouyang et al., 2022), where feedback is collected based on comparisons between outputs. However, instead of relying on real human feedback, we employed a strong large language model (LLM) model as a judge to simulate human interactions with the chatbot and provide feedback. For real-world online marketplace chatbots like the Amazon Rufus (Amazon, 2024), human feedback is typically given as a rating at the end of the interaction. However, human users tend to compare their current experience with previous ones when assigning ratings. To capture this behavior, we simulate users comparing services from two different stores and selecting their preferred option, rather than rating each scenario in isolation. This closely aligns with the preference-based data collection method used in prior work (Stiennon et al., 2020; Ouyang et al., 2022), where users provide feedback by comparing responses instead of giving individual ratings.
Decision-making simulation. While collecting the preference data, our simulated human (strong model) takes on three roles: interacting with the chatbot, making decisions, and providing feedback. To ensure accurate decision-making and feedback, we adapted the approach in introspective planning (Liang et al., ). First, we frame the decision-making problem as a multiple-choice question with four options: (A) Buy option A, (B) Buy option B, (C) Buy option C, or (D) Do not buy anything. We then ask the LLMs to perform Chain-of-Thought reasoning (Wei et al., 2022), querying the next token probabilities to select the best option from ${A,B,C,D}$. This approach can reduce the language agent’s uncertainty. We apply a similar method for comparing services between two stores.
Dataset Details. In our experiments, we used both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) as the AI assistants, and Llama-3.1-70B (Dubey et al., 2024) as the simulated human to interact with the AI assistant and provide feedback. We collected 11,000 preference data points for each AI assistant model, with 10,000 used for training and 1,000 for validation. We also generated a test set of 1,200 examples to evaluate performance across different customer scenarios.
4.2 Experiment Setup
Environment Details. In each of our simulated marketplace scenarios there are 10 candidate items, each characterized by 8 features and a price. Each feature can be categorized in two ways: (1) The item either has or lacks a specific feature (e.g., a TV with HDR vs. without HDR), and (2) The feature can vary in types (e.g., 8K resolution vs. 4K resolution). While in most cases the chatbot has access to this information, there are instances where it is uncertain about a particular feature (e.g., resolution not specified). We will examine these scenarios and investigate when and how the AI acts deceptively.
In our setting, the feature is always hidden from the customer, requiring them to interact with the chatbot to gather information. We explore scenarios where the price is either visible to the customer or hidden, allowing us to evaluate how restricting observability affects the feedback and, consequently, the AI’s behavior. We also consider scenarios when the customer prioritizes price by adding a constraint regarding their price requirements in the prompt.
Metrics. We use two primary metrics: true utility and satisfaction rating. The true utility metric $U$ reflects both the customer’s requirements and the item they purchase. We define $U$ as follows: if the customer makes no purchase, the utility is $U=0$. If the purchased item lacks the required feature, $U=-1$. If the item contains the required feature and the customer has no price constraints, $U=1$. When price is a priority and the item contains the required feature, the utility is defined as the ratio of the price of the cheapest item with the feature to the price the customer actually paid.
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale ranging from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Fig. 4 and Fig. 5, these ratings were normalized to a scale between -1 and 1, which ensure that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. Furthermore, we quantified two metrics in the human study: regret rate, which measures how often users regret their decisions, and hallucination rate, which measures how truthful the language model is.
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data. The language model then interacted with the environment by generating responses and receiving reward signals from this reward model. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tuned the model iteratively to maximize these rewards. For the offline approach, we experimented with Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without the need for an explicit reward model. We apply LoRA fine-tuning (Hu et al., 2021) for both algorithms to efficiently update model parameters. Further details of these methods are included in Appendix B.
Figure 4: Results on Llama-2-7b trained with PPO. Left: Demonstrates the Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Shows how partial hindsight mitigate the misalignment. Right: Shows the alignment achieved with oracle hindsight.
Figure 5: Results on Llama-2-7b trained with DPO. Left: Demonstrates the Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Shows how partial hindsight mitigate the misalignment. Right: Shows the alignment achieved with oracle hindsight.
5 Simulation Results
Misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe significant misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4 and 5). While the satisfaction rating steadily increases, indicating that the language model is learning to deliver responses that receive higher immediate user approval, the true utility shows a sharp decline. This suggests that while the chatbot’s responses may appear more polished or helpful in the short term, they are in fact becoming less aligned with the user’s true needs or long-term goals. As a result, while users may initially perceive the chatbot’s responses as helpful, they are frequently misled and ultimately dissatisfied with their final outcomes. This highlights a fundamental flaw in using standard RLHF with immediate feedback, as it risks optimizing for superficial satisfaction at the expense of true utility.
Metric DPO PPO
IF PHS OHS IF PHS OHS
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.33_{\pm 0.036}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.31_{\pm 0.024}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $0.23_{\pm 0.026}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $0.24_{\pm 0.031}$
Table 1: Comparison of performance metrics (Rating and True Utility) across models trained with DPO and PPO under three feedback conditions: Immediate Feedback (IF), Partial Hindsight Simulation (PHS), and Oracle Hindsight Simulation (OHS). Ratings are higher when trained with immediate feedback but lead to lower real utility, indicating potential misalignment between perceived satisfaction and actual utility. Hindsight simulations significantly improve the true utility.
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), relying on immediate feedback leads to a steady decline in real utility, ultimately resulting in negative overall utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them. The qualitative results shown in Fig. 3 further support our claim. When the AI assistant is trained on immediate feedback, it deceptively claims that both Options A and C meet the requirements of the (simulated) customer for 8K resolution, though neither actually does. In contrast, training with partial hindsight leads to truthful responses, acknowledging that none of the options meet the 8K resolution requirement. This shows that while traditional RLHF with immediate feedback may cause misalignment, hindsight simulation mitigates this issue, improving the overall helpfulness and honesty of language agents.
6 Human Study
Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix D.2.
Figure 6: The policy trained using the proposed RLHS outperforms that of RLHF in both true utility (left) and hindsight rating (right). In both plots, each point represents the mean ratio for a scenario, with lines indicating the standard deviation. The identity line is plotted in dashed grey.
Pipeline for evaluating model performance. The first and second experiments follow the same pipeline but differ in the models used—one is trained with immediate feedback, and the other with partial hindsight simulation—allowing us to compare their performance (Goal 1). Initially, participants are shown a list of available items in a store with hidden features. We specify their requirements for the item (e.g., “must have 8K resolution”). Participants interact with the chatbot to gather information about the products. At each step, they can choose one of the following actions: “ask about the desired feature,” “ask about the price”, or “ready to make a decision”. Pre-generated responses are provided for inquiries. In the second round of interaction, participants may ask about the information they didn’t request in the first round. At any point, participants can choose “ready to make a decision”, at which time they must decide whether to make a purchase decision or opt not to buy. After making their decision, they provide an immediate satisfaction rating.
Hindsight information is then introduced. Buyers learn whether the item meets their requirements (e.g., whether the item has the desired feature) while non-buyers receive no additional information. Participants then provide a second satisfaction rating, referred to as the hindsight rating, which evaluates their long-term satisfaction after considering the hindsight information. This step allows us to assess the impact of hindsight information on user satisfaction (Goal 2). Finally, buyers may keep or return the item, enabling us to quantify the regret rate.
Statistical Hypothesis Testing. We conducted experiments to test four hypotheses, using one-tailed and standard t-tests for the first three hypotheses (Fisher, 1970), and Pearson’s correlation coefficient for the fourth (Sedgwick, 2012). The one-tailed t-test used for Hypotheses 1, 2, and 3 is outlined below. The null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{1}$) are defined as:
[…]
Here, $\mu_{1}$ and $\mu_{2}$ represent the mean satisfaction for Group 1 and Group 2, respectively. The two-tailed t-test follows a similar format but tests for any significant difference between the group means.
Hypothesis 1: Models trained with RLHS lead to a higher long-term user satisfaction rate and a lower regret rate than those trained with RLHF using immediate feedback.
We evaluated hindsight ratings for two models: Group 1 (RLHS) and Group 2 (RLHF). The hypothesis test resulted in $p=4\times 10^{-8}$, well below the significance threshold of 0.001. When reversing the groups for regret rates, the test yielded $p=5\times 10^{-5}$ again below 0.001.
Hypothesis 2: Models trained with RLHF using immediate feedback often experience a notable decline in user satisfaction once future outcomes are revealed, and RLHS mitigates this decline.
Group 1 consisted of users interacting with RLHF without hindsight feedback, and Group 2 received hindsight feedback. The hypothesis test gave $p=4\times 10^{-9}$, confirming a significant decline in user satisfaction. To demonstrate that RLHS mitigates this decline, we ran a two-tailed t-test comparing immediate and hindsight ratings, and find that this decline is likely no longer present ($p=0.90$).
Hypothesis 3: RLHS leads to significantly higher true utility than RLHF.
We assessed the objective performance of the two models by comparing true utility scores for Group 1 (RLHS) and Group 2 (RLHF). The hypothesis test yielded $p=4\times 10^{-8}$.
Hypothesis 4: Models trained with RLHS are more truthful, presenting a strong correlation between their high immediate user satisfaction rate (subjective) and high true utility (objective).
To evaluate the correlation, we used Pearson’s correlation coefficient and tested whether this coefficient was significantly different from zero. The null hypothesis ($H_{0}$) assumed no correlation (i.e., $r=0$) while the alternative hypothesis ($H_{1}$) assumed a non-zero correlation. The test found a significant correlation between immediate ratings and true utility for RLHS ($p=5\times 10^{-4}$), while no significant correlation was observed for RLHF ($p=0.47$).
Model Immediate rating Hindsight rating True utility Regret rate
RLHF $3.74_{\pm 0.94}$ $2.65_{\pm 1.55}$ $-0.16_{\pm 0.87}$ $0.64_{\pm 0.48}$
RLHS $3.69_{\pm 1.05}$ $3.71_{\pm 1.10}$ $0.43_{\pm 0.60}$ $0.23_{\pm 0.42}$
Table 2: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS outperforms in hindsight rating, true utility, and regret rate, indicating better long-term alignment with user preferences and reduced regret.
Analysis. Statistical significance tests verified Hypotheses 1–4. As shown in Table 2, RLHS significantly outperformed RLHF by achieving higher hindsight satisfaction scores (3.71 vs. 2.65), higher true utility (0.43 vs. -0.16), and lower regret rates (0.23 vs. 0.64). These results demonstrate the alignment and performance advantages of RLHS over RLHF. We also visualize the utility and rating for each scenario in Fig. 6. RLHS consistently achieves higher true utility and hindsight ratings compared to RLHF in most scenarios, demonstrating its superior alignment and performance. Additionally, we analyzed the hallucination rate across 10 scenarios. RLHS reduced the hallucination rate from $80\%$ (RLHF) to $0\%$, demonstrating models’ enhanced truthfulness using our approach.
7 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Wang et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, producing data used to train a fixed reward model. This reward model is then applied in reinforcement learning, a process referred to as RLAIF. The technique of using LLM-as-a-Judge has become a standard method for evaluating model outputs (Dubois et al., 2024; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023a). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
Challenges of Learning from Human Feedback. Learning from human feedback presents challenges (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes due to limited time (Chmielewski & Kucker, 2020), incomplete information (Casper et al., 2023; Lang et al., 2024), lack of expertise (Daniels-Koch & Freedman, 2022) or cognitive biases (Pandey et al., 2022). Evaluators may also have conflicting preferences (Bakker et al., 2022). Modeling human preferences is difficult (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022), with models being prone to overoptimization (Gao et al., 2023). Due to the imperfect nature of human judgment, we argue that relying on immediate feedback, as is common in current RLHF pipelines, can lead to misalignment. In this work, we propose a hindsight simulation approach that aims to reduce human uncertainty and foster more truthful feedback, thereby mitigating these alignment challenges.
8 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with future outcome information. Using both theoretical proofs and realistic human experiments, we demonstrate that RLHS can significantly improve utility compared to existing RLHF pipelines that rely on only immediate feedback, while maintaining a high user satisfaction rate throughout the human–AI interactive process. While our study focused on simulated hindsight with an application to marketplace chatbots, future work should explore incorporating hindsight in RLHF for additional real-world applications with real human evaluators. We also see open opportunities to equip RLHS with other feedback modalities, such as visual cues, which could enrich the feedback process and improve alignment.
Appendix A Additional Quantitative Results
Figure 7: Results on Llama-3-8b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Right: Partial hindsight mitigate the misalignment.
Figure 8: Results on Llama-3-8b trained with DPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Right: Partial hindsight mitigate the misalignment.
(a) PPO training result
(b) DPO training result
Figure 9: Likert scale satisfaction ratings for Llama-3-8b. The comparison includes ratings for Immediate Feedback (grey), Partial Hindsight (orange).
(a) PPO training result
(b) DPO training result
Figure 10: Likert scale satisfaction ratings for Llama-2-7b. The comparison includes ratings for Immediate Feedback (grey), Partial Hindsight (orange), and Oracle Hindsight (green).
(a) Immediate feedback
(b) Partial hindsight
Figure 11: Histograms of Likert ratings for Llama-2-7b trained with PPO using immediate feedback (a) and partial hindsight (b). The model trained with immediate feedback achieves high ratings (predominantly 5), but has a negative true utility (-0.71), indicating significant misalignment. In contrast, the model trained with partial hindsight maintains high ratings while achieving high true utility (+0.18), demonstrating better alignment between user ratings and true utility.
Analysis: We provided additional experimental results on Llama-3-8b using PPO and DPO in Fig. 7 and Fig. 8. The results further justifies our claim on misalignment and the effectiveness of hindsight to mitigate the misalignment. We also provided the Likert scale satisfaction ratings for both Llama-2-7b and Llama-3-8b in Fig. 9 and Fig. 10 and conducted additional analysis of the distribution of the ratings in Fig. 11. We observed that models trained with immediate feedback achieve very high satisfaction ratings (predominantly 5), as illustrated in the histogram in Fig. 11(a). However, this comes at the expense of true utility (-0.71), which remains negative and underscores the misalignment issue between satisfaction and true utility. Training with hindsight feedback still maintains a high satisfaction rating while significantly improving true utility, achieving positive values (+0.18), as shown in Fig. 11(b). This indicates that partial hindsight mitigates the misalignment, resulting in more truthful model performance.
Metric DPO PPO SimPO
IF PHS IF PHS IF PHS
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.94_{\pm 0.032}$ $0.37_{\pm 0.028}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $-0.49_{\pm 0.044}$ $0.16_{\pm 0.032}$
Table 3: Performance comparison of DPO, PPO, and SimPO models under Immediate Feedback (IF) and Partial Hindsight Simulation (PHS). Average satisfaction ratings and true utility (with standard deviations) are shown. SimPO results are included for comparison between online (PPO) and offline (DPO, SimPO) RLHF approaches.
Comparison between online and offline fine-tuning. We ran both t-test and two-way ANOVA to better understand emergent misalignment and the effectiveness of mitigation through hindsight simulation under online and offline fine-tuning schemes. Results show that PPO with immediate feedback yields significantly lower true utility for the user than DPO ($p=1.1\times 10^{-4}$ in t-test). In addition, considering the difference between the (normalized) user rating and true utility, we find that immediate feedback in online RLHF using PPO introduces a larger misalignment gap than offline RLHF using DPO ($p=6.7\times 10^{-5}$ in t-test). Incorporating partial hindsight helps mitigate this misalignment gap across online and offline fine-tuning ($p=3.1\times 10^{-116}$ in two-way ANOVA test). We also compared online PPO with offline SimPO (Meng et al., 2024) and found that PPO introduces a larger misalignment gap than SimPO ($p=8.2\times 10^{-5}$ in t-test), with partial hindsight significantly reducing misalignment in SimPO as well ($p=5\times 10^{-56}$ in t-test).
Appendix B Training algorithms.
The initial stage of alignment involves Supervised Fine-Tuning (SFT), where the pre-trained model is adapted to mimic high-quality demonstration data, such as dialogues and summaries. To enhance alignment of the SFT model $\pi_{\theta}$ with human preferences, previous studies (Ziegler et al., 2019; Ouyang et al., 2022) have implemented the Reinforcement Learning from Human Feedback (RLHF) technique. This approach optimizes the following objective:
$J_{r}(\pi_{\theta})=\mathbb{E}_{\mathbf{x}\sim p_{\text{data}},\mathbf{y}\sim\pi_{\theta}}\left[r(\mathbf{x},\mathbf{y})-\beta\log\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})}\right],$ (4)
where $r(\mathbf{x},\mathbf{y})$ is the reward function reflecting human preferences, $\pi_{\theta}$ is a policy model, and $\pi_{\text{ref}}$ is a reference policy used for regularizing $\pi_{\theta}$ with Kullback–Leibler divergence. The term $\beta$ is a regularization parameter to control the degree of regularization.
Online preference optimization. When the reward $r$ is unknown, a reward model $r_{\phi}$ is derived from human-labeled data. This dataset consists of pairs $(x,y_{w},y_{l})$, with $y_{w}$ and $y_{l}$ designated as the preferred and less preferred responses by human evaluators respectively. The preference likelihood, as per the Bradley-Terry model (Bradley & Terry, 1952), is given by:
$\mathbb{P}(y_{w}>y_{l}\mid x)=\frac{\exp(r_{\phi}(x,y_{w}))}{\exp(r_{\phi}(x,y_{w}))+\exp(r_{\phi}(x,y_{l}))}$
To optimize $r_{\phi}$, we minimize the negative log-likelihood of this model:
$L_{R}(r_{\phi})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$
Once $r_{\phi}$ is fine-tuned, it substitutes the initial reward function $r$ and is integrated directly into the reinforcement learning framework, enhancing the model’s performance through explicit optimization via Proximal Policy Optimization (PPO) (Schulman et al., 2017):
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim p_{\nu}}\left[r_{\phi}(x,y)-\beta D_{KL}(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x))\right]$
Here, $\beta$ adjusts the deviation from the base reference policy $\pi_{\text{ref}}$, ensuring the model adheres closely to desired behaviors.
Offline preference optimization. We experimented with Direct Preference Optimization (DPO), which aligns language models with human preferences without the need for an explicit reward model. DPO reparameterizes the reward function $r$ using the following expression:
$r(\mathbf{x},\mathbf{y})=\beta\log\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})}+\beta\log Z(\mathbf{x})$ (5)
where $Z(\mathbf{x})$ is the partition function. The objective for DPO then becomes:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(\mathbf{y}_{w}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}_{w}|\mathbf{x})}-\beta\log\frac{\pi_{\theta}(\mathbf{y}_{l}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}_{l}|\mathbf{x})}\right)\right],$ (6)
where $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ are preference pairs consisting of the prompt, the winning response, and the losing response from the preference dataset $\mathcal{D}$. This formulation allows DPO to optimize directly based on preferences without requiring a reward model.
Appendix C Prompts
In this section, we provide example prompts that we used to generate feedback from AI.
Table 4: Prompt for generating feedback from AI.
You are a customer seeking to make a purchase, and you visit two different online stores, both of which have a chatbot that you can consult to help you choose.
Both stores contain three options on {item}:
Option A: {price_A}
Option B: {price_B}
Option C: {price_C}
Your absolute requirement is a {item} that must include the following features: {requirement}.
Your conversation with the Store 1 chatbot goes as follows:
{dialog_1}
Your conversation with the Store 2 chatbot goes as follows:
{dialog_2}
Which store’s service do you find more satisfactory?
Table 5: Prompt for asking the customer to make decision.
Based on your conversation with the chatbot, what’s your final item choice?
A) Buy Option A
B) Buy Option B
C) Buy Option C
D) Do not buy anything
Please answer with a capital letter (A, B, C, or D) and explain why.
Table 6: Prompt for asking the customer’s satisfaction.
How satisfied are you with the service? Please provide a rating from 1 to 5, based on the following criteria:
1: Very Dissatisfied
   - The chatbot provided misleading or incorrect information.
   - The chatbot’s responses were irrelevant or off-topic.
   - The interaction was frustrating, and my questions were not addressed.
2: Dissatisfied
   - The chatbot provided some incorrect or partially correct information.
   - The chatbot’s responses were occasionally irrelevant or not helpful.
   - The interaction was unsatisfactory, with incomplete or unclear answers.
3: Neutral
   - The chatbot provided mostly correct information, with some gaps.
   - The chatbot’s responses were somewhat helpful but not comprehensive.
   - The interaction was average, neither particularly good nor bad.
4: Satisfied
   - The chatbot provided accurate and helpful information.
   - The chatbot’s responses were relevant and mostly thorough.
   - The interaction was smooth, addressing most of my questions effectively.
5: Very Satisfied
   - The chatbot provided clear, accurate, and highly helpful information.
   - The chatbot’s responses were comprehensive and directly addressed my needs.
   - The interaction was excellent, resolving my queries effectively and efficiently.
Please rate your experience from 1 to 5 based on the above criteria and explain the reasons for your rating.
Appendix D Human Study Details
D.1 User Interface
In this subsection, we display the interface used in our human study.
Figure 12: Example of user interaction interface for our main human experiments studying the misalignment of RLHF and the effecitveness of RLHS.
Figure 13: Example of user interaction interface for additional human experiments assessing the alignment of LLM actions and feedback with those of humans.
D.2 Participants and data collection
The human subjects were chosen from a high quality Prolific participant pool, where participants were pre-screened to have an approval rate of 95-100 over at least 100 previous submissions. Participants were located in the USA. To assign subjects to experimental conditions, we used random assignment, and each participant was only assigned to one shopping scenario (either one purchasing decision or comparing between two AI shopping assistants). As a negative experience could bias participants’ perceptions of AI chatbots, we ensured that they were not able to retake the study.
The expected duration of the study was 5 minutes, and actually completed the study at a median time of 4:54. Subjects were compensated \$1.10 for their participation, resulting in a hourly wage of \$13.47/hour, which was substantially higher than minimum wage. In addition to participant satisfaction ratings or preferences, participants were asked to provide a brief 2-sentence explanation to explain their ratings or preferences. We manually reviewed these explanations for all participants, and participants that did not provide a reasonable 2-sentence explanation had their data removed from the study. We also removed participants that finished the study in an unreasonably short time ($<$1:30 out of the estimated 5 minutes). Other than this, no data was removed.
This study received IRB approval at Princeton University with the record number 10859.
D.3 Additional human study
We conducted an additional human study to assess how closely the feedback and actions of our AI proxy (Llama-3.1-70B) align with those of human participants. In the study, participants interacted with chatbots from two different stores, taking actions such as purchasing items or leaving the store based on the conversations. After engaging with both stores, participants were asked to choose which store they preferred. We randomly selected 10 scenarios from our training set, with 30 different participants evaluating each scenario. To determine the human preference for each scenario, we employed majority voting. This method was used to ensure that the aggregated choice reflected the consensus among participants, minimizing the impact of individual variability and providing a more robust measure of overall preference. Our analysis revealed that the matching accuracy between LLM-generated feedback and human feedback reached $100\%$. Furthermore, the actions taken by the LLM matched those of human participants with $95\%$ accuracy. These findings suggest that our simulated feedback and actions align strongly with real human behavior.
Appendix E Discussion
E.1 Related Work
Statement of Contributions. Our key insight is that the true value of AI outputs lies in their downstream consequences, especially in how they influence real-world human behavior. While the importance of long-term outcomes is a fundamental aspect of dynamic decision theory, our work is the first to address this within the context of RLHF by (1) exploring the negative effects of learning from immediate human feedback based on foresight, and (2) proposing a general mitigation strategy that evaluates real-world downstream harm caused by inaccurate information.
Comparison with Related Work: One of the recent works cited in comparison is by Lang et al. (2024), which focuses on the problem of partial observability. This is distinct from the problem of human misprediction we address. In their setting, user utility is confined to the immediate time frame of the interaction and does not consider the long-term repercussions on the user’s behavior or well-being after the interaction concludes. Their analysis primarily highlights scenarios where an AI system is incentivized to withhold information to avoid negative feedback scores but does not delve into the real-world impact such deception has on user utility. In contrast, our approach specifically examines the human user’s decision-making process after interacting with the AI system, emphasizing how misalignment or deceptive behavior directly affects their realized utility.
Recent studies have investigated sycophantic behavior in language models (Sharma et al., 2023; Wei et al., 2023; Perez et al., 2022), where the models are optimized to generate responses that align with user beliefs rather than the truth. Our empirical results also reveal such tendencies. In this paper, we analyze the underlying factors contributing to this behavior and demonstrate how incorporating hindsight can be effective in preventing sycophancy.
Theoretical Contributions: We extend the RLHF formulation by mathematically capturing this dynamic interplay between AI and human decision-making, something that has not been explored in prior work, including Lang et al. (2024) Our theoretical analysis not only highlights why deceptive behavior is problematic but also quantifies its repercussions by modeling the "closed-loop" evolution of the sociotechnical system formed by the human and the AI.
Mitigation Strategies: Importantly, we propose and evaluate a novel mitigation method: Reinforcement Learning from Hindsight Simulation (RLHS) that significantly reduces misalignment and deceptive behavior. While Lang et al. (2024) note that deception is undesirable, they do not provide solutions or a theoretical basis for understanding its downstream damage. Our work, therefore, not only identifies and analyzes the issue but also offers a practical, effective mitigation strategy. Additionally, our partial hindsight approach still operates within a partially observable setting. The minimal difference between partial and oracle hindsight suggests that the fundamental issue in the class of misalignment we study is not primarily linked to partial observability, but rather to the human misprediction of the downstream consequences.
E.2 Broader Impact
Human evaluators do not always know the full truth when providing feedback. Without explicit information about its future consequences, evaluators must implicitly estimate them during their assessment. This limitation poses significant challenges for real-world applications of AI, particularly within the RLHF framework we studied. In the following sections, we discussed these limitations and how our proposed hindsight feedback approach can help overcome them to enhance AI alignment.
Limited Access in Real-World Applications: In real-world scenarios, users and human labelers frequently interact with black-box or closed-prompt AI systems where internal prompts and decision-making processes remain opaque. Notable examples include commercial systems such as OpenAI’s ChatGPT and Amazon’s Rufus. Our proposed techniques (hindsight feedback), and the experimental settings we used can be applied directly to these systems where full internal access is unavailable. In such cases, assessing the consistency of responses alone is insufficient, as external context might not capture the complete implications of an AI’s output. Hindsight feedback allows evaluators to provide more reliable feedback by considering outcomes, improving alignment in these constrained settings.
Limitations of Human Judgment and Information Access: Even when human evaluators have full access to models and their prompts (e.g., in open-source systems), perfect judgment is not guaranteed. Evaluators may miss deeper implications or fail to predict the long-term impact of responses, whether due to lack of expertise or cognitive limitations. These challenges are relevant to both open-source and closed-source models. Below, we outline two practical examples illustrating these limitations and how hindsight feedback can address them:
Code Generation Scenario: Imagine a user asking a language model for code to fit a polynomial curve to a set of data points. One solution may fit the data perfectly, while another shows some deviation. A human evaluator might prefer the model with the perfect fit, not realizing that it overfits and performs poorly on new data. Immediate feedback in this case could lead to misalignment, as it prioritizes surface-level satisfaction over long-term utility. By providing feedback after testing the code on new data (hindsight), evaluators can offer more informed input, reducing misalignment. Hindsight simulation can automate part of this process by allowing models to test outcomes on unseen data and report the results to human evaluators for better feedback. One extra benefit of hindsight simulation is that humans do not need to be domain experts to provide truthful feedback.
AI4Science Proof Construction: When constructing mathematical proofs for scientific problems, model may generate results that are correct only under conditions or assumptions specified by the user. Human evaluators, constrained by time or limited expertise, may overlook these limitations during evaluating the model, eventually causing the model to overfit to a restricted set of problems and unable to tackle scientific problems in general settings. On the other hand, hindsight simulation generates a diverse set of scenarios, including, e.g., edge cases, under which the model is required to validate its proof. This allows the human evaluator to assess the model performance based on its ability to generalize beyond the immediate problem.
Algorithm 1 Human Feedback Loop for RLHS
1:  Step 0: Initialization
2:  ${s}_{0},{z}^{\textit{H}}_{0},{\theta}^{\textit{H}},{o}_{0}^{H}\leftarrow\text{sample\_initial\_conditions}(\mathcal{S},\mathcal{Z}^{\textit{H}},\Theta^{\textit{H}})$
3:
4:  Step 1: AI Prompt Sampling
5:  ${s}_{0}^{\textit{AI}},{o}_{0}^{\textit{AI}}\leftarrow\text{sample\_AI\_prompt}(\mathcal{Z}^{\textit{AI}},\mathcal{O}^{\textit{AI}})$
6:
7:  Step 2: AI Policy Evaluation
8:  Query the AI policy for an action: ${o}^{H}_{1}:=a^{\textit{AI}}_{0}\sim{\pi}^{\textit{AI}}(\cdot\mid{s}_{0},{z}^{\textit{H}}_{0})$
9:
10:  Step 3: Hindsight
11:  for $t=1$ to $T+N$ do
12:     Sample action: $a_{t}\leftarrow\text{sample\_action}({\pi}^{\textit{AI}})$
13:     ${s}_{t+1},{o}_{t+1}^{H}\leftarrow f({s}_{t},a_{t},{o}_{t}^{\textit{H}})$
14:  end for
15:
16:  Step 4: Query Feedback
17:  Query human feedback on the AI policy: $\hat{{U}}^{H}(\pi^{AI})\leftarrow\text{query\_human\_feedback}(\pi^{AI})$
18:
19:  Output or Process Feedback
[…]
Notation Marketplace AI4Science Proof Construction
${s}_{0}$ Initial store inventory Initial problem instance
$\Theta_{0}$ Customer’s desired features User goals/preferences (succeed at future problem instances)
${s}_{0}^{H}$ Background knowledge Prior knowledge about the problem
${o}_{0}^{H}$ Available public stock information Initial problem setup
$z_{0}^{\text{AI}}$ AI system’s internal information Initial problem setup
${o}_{0}^{\text{AI}}$ All detailed stock information Initial problem setup
$a_{t}^{H}$ Customer’s follow-up question or purchase decision User’s input or solution attempt
${s}_{t+1}$ Product arrival Next problem instance
${o}_{t+1}^{H}$ Revealed product features Validation or correctness check
$\hat{{U}}^{H}$ Satisfaction with the service Satisfaction with the solution
Table 7: RLHS notations for Marketplace and AI4Science Proof Construction
Appendix F Additional Qualitative Results
In this section, we provide additional results comparing the qualitative differences between the outputs of policies trained with RLHF and RLHS. We also show a failure case here.
Figure 14: Qualitative results for Llama-2-7b trained with DPO using immediate feedback versus partial hindsight. The model trained with immediate feedback falsely claims that Option B is most affordable with 8K resolution, which is incorrect. In contrast, the model trained with partial hindsight truthfully states that option C is the most affordable option that includes 8K resolution.
Figure 15: Qualitative results for Llama-3-8b trained with DPO using immediate feedback versus partial hindsight. The model trained with immediate feedback falsely claims that Option C can play 3D movies, which is incorrect. In contrast, the model trained with partial hindsight accurately states that Option C’s 3D capability is not specified, and recommends Option B, the cheapest option that includes 3D capability.
Figure 16: Failure case for Llama-2-7b trained with DPO using partial hindsight. The model trained with immediate feedback deceives about specific features, while the model trained with partial hindsight withholds some information. This reveals shortcomings of partial hindsight, as it does not have observations for all other items. Consequently, it might still encourage the agent to deceive about the price or conceal price information.