### v2

# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

Abstract
While Reinforcement Learning from Human Feedback (RLHF) has shown promise in aligning generative AI, we present empirical evidence that it can also cause severe, systematic misalignment. We hypothesize that this stems from evaluator feedback depending on downstream outcome predictions (foresight) that can be influenced by the AI’s output, inducing Goodhart’s law dynamics. Conversely, our theoretical analysis shows that conditioning evaluator feedback on downstream observations (hindsight) inhibits this effect by decoupling the alignment signal from potentially compromised predictions—crucially, the result holds even if the observed outcomes are sampled from the AI’s own world model. Building on this insight, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which presents plausible simulated outcomes to evaluators before eliciting feedback. We demonstrate RLHS on online (PPO) and offline (DPO) large language model fine-tuning, obtaining superior alignment over RLHF in controlled consultancy-type experiments and user studies. We evaluate post-hoc on the TruthfulQA benchmark and find that, even after single-task fine-tuning, both RLHF misalignment and RLHS alignment carry over to substantially different settings.
Machine Learning, ICML
1 Introduction
Aligning artificial intelligence (AI) systems with human values and goals is crucial to ensuring their behavior is helpful, honest, and trustworthy. Eliciting human feedback is a widely-used alignment strategy (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it relies predominantly on immediate assessments of isolated interactions, which may not accurately account for their downstream outcomes. Inaccurate user or evaluator feedback can misguide the model’s behavior and undermine the alignment process. For example, human evaluators could misjudge an interaction on the spot, due to limited knowledge (Casper et al., 2023; Lang et al., 2024) or cognitive resources  (Pandey et al., 2022; Chmielewski & Kucker, 2020), leading to incomplete or misinformed feedback. A recent study has theorized that partial observability of an AI assistant’s task execution during human–AI interaction can lead RLHF to learn deceptive behaviors (Lang et al., 2024).
Figure 1: RLHF can incentivize AI systems to provide inaccurate or deceptive information to their users, prioritizing positive on-the-spot feedback and neglecting long-term consequences. For example, a customer may prefer to hear good news while shopping but will ultimately be disappointed (and objectively worse off) if stuck with an ill-informed purchase. The proposed RLHS scheme removes this incentive by simulating the interaction’s downstream outcome before eliciting feedback. The resulting fine-tuned AI models show superior alignment with users’ underlying utility.
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. We contend that rewarding an AI system to improve users or evaluators’ in-the-moment assessments of interactions creates a pernicious Goodhart’s law dynamic: such an incentive will favor outputs that induce unrealistically optimistic expectations in users—while at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded outcomes.
We provide substantial empirical evidence that indeed this phenomenon can arise even in simple settings: we find that immediate human feedback elicited at the end of the human–AI interaction frequently misrepresents true utility in consultancy-type interactions, and, when used as a proxy for it in RLHF fine-tuning, it systematically drives misalignment with human goals (Fig. 1, top). Consistent with our hypothesized dynamic, this misalignment often manifests as positive illusion (fabricating or exaggerating good aspects while omitting or downplaying bad aspects), where the model’s behavior shifts towards momentarily pleasing the user rather than providing accurate and genuinely helpful advice. This consistently leads users to make ill-informed decisions whose poor downstream outcomes contrast starkly with their high satisfaction rating at the end of the interaction.
To address these open challenges, we propose to leverage hindsight as a simple but effective misalignment mitigation mechanism, in which evaluators experience the downstream outcomes of an interaction before being asked for feedback on the model. We provide both theoretical analysis and extensive empirical studies to show the efficacy of hindsight in significantly reducing misalignment of RLHF-trained models. To circumvent the material and ethical difficulties in exposing real people to real consequences, we introduce a novel alignment algorithm called Reinforcement Learning from Hindsight Simulation (RLHS), an alternative to RLHF that uses an AI world model during training to simulate likely human decisions and their downstream outcomes. Despite lacking any privileged information on real outcomes, the method effectively relieves evaluators of the need to predict the future consequences of an AI output—thus removing the AI incentive to influence these predictions.
Our key finding is that equipping evaluators with the benefit of (simulated) hindsight—and thereby relieving them of the burden of prediction—significantly reduces model misalignment after fine-tuning. Intuitively, even if the world model generating outcomes contains inaccuracies, these are not dependent on the AI output, and therefore the AI agent has no incentive to manipulate them. We evaluate hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and find that it greatly improves alignment in both training paradigms. We validate these results through an online user study, in which RLHS consistently improved the objective utility and subjective satisfaction of our human participants, despite being trained with only simulated hindsight feedback. Our comparative findings demonstrate that RLHS significantly outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for human feedback, and has been shown to produce results similar to RLHF (Bai et al., 2022b; Lee et al., 2023). Finally, a post-hoc evaluation on the TruthfulQA benchmark indicates that, even after single-task fine-tuning, both RLHF misalignment and RLHS alignment can persist in substantially different settings.
2 Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We model the a problem as a partially observable Markov decision process (POMDP) defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$) encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
AI-Assisted Human Decision-Making. When the human consults an AI system (e.g., a FM) to help with their decision problem, we may augment the above problem with the human–AI interaction. The resulting Assisted POMDP is a tuple $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}=(\mathcal{S},\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\mathcal{O}^{\textit{H}},\mathcal{O}^{\textit{AI}},\mathcal{T},{O}^{\textit{H}},{O}^{\textit{AI}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ and $\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ are the sets of interactive actions available to the human and AI system, $\mathcal{O}^{\textit{AI}}$ is the AI’s observation space, and ${O}^{\textit{AI}}$ is the AI’s observation map ${O}^{\textit{AI}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{AI}})$. In this model, the AI takes an advisory role: it can respond to a human’s interactive action ${a}_{\rightleftharpoons}^{\textit{H}}\in\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ (e.g., a query through a chat interface) with its own ${a}^{\textit{AI}}_{\rightleftharpoons}\in\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ (e.g., a generated text or multimedia output). After one or multiple rounds of such interactions, the human may take a physical action ${a}^{\textit{H}}\in\mathcal{A}^{\textit{H}}$ to affect the evolution of the world state ${s}$. This Assisted POMDP is a special case of a partially observable stochastic game (POSG) (Hansen et al., 2004). In such interactions, the AI’s goal is to influence the human’s internal state ${z}^{\textit{H}}$ towards maximizing the rewards $r({s},{a}^{\textit{H}};{\theta}^{\textit{H}})$ accrued over time. This, however, is made challenging by the AI’s fundamental uncertainty about the human’s preferences ${\theta}^{\textit{H}}$.
Reinforcement Learning from Human Feedback (RLHF). RLHF aims to learn the human’s preferences ${\theta}^{\textit{H}}$ from human feedback data, which typically involves three key steps. In Step 1, the human is asked to provide feedback on some state sequences ${\mathbf{{s}}}=({s}_{0},{s}_{1},\ldots,{s}_{k})$ (e.g., a human–AI dialogue), with ${s}_{t}\in\mathcal{S},~{}\forall t=0,1,\ldots,k$. For example, in binary comparison (Christiano et al., 2017), assuming human is a Boltzmann-rational decision maker (Luce, 1959), the probability that the human prefers ${\mathbf{{s}}}$ over ${\mathbf{{s}}}^{\prime}$ is $P^{r}_{k}({\mathbf{{s}}}\succ{\mathbf{{s}}}^{\prime})=\sigma(\beta(R_{k}({\mathbf{{s}}})-R_{k}({\mathbf{{s}}}^{\prime})))$, where $\sigma(\cdot)$ is the sigmoid function, $\beta>0$ is the inverse temperature parameter, and $R_{k}({\mathbf{{s}}})=\sum_{t=0}^{T}\gamma^{t}r({s}_{t})$ is the return received by state sequence ${\mathbf{{s}}}$. Step 2 is to fit a reward function $\hat{r}$ based on a dataset containing state sequences paired with human feedback, hoping that $\hat{r}$ will resemble $r$ as closely as possible. Step 3 is to compute an AI policy $\hat{{\pi}}:\mathcal{S}\rightarrow\Delta(\mathcal{A}^{\textit{AI}}_{\rightleftharpoons})$ that maximizes the return based on the estimated reward $\hat{r}$, i.e., $\hat{{\pi}}=\operatorname*{arg\,max}_{\pi}{U}_{k}({\pi})$, where ${U}_{k}({\pi}):=\operatorname*{\mathbb{E}}_{{\mathbf{{s}}}\sim p^{\pi}}[\hat{R}_{k}({\mathbf{{s}}})]$ is the expected utility of ${\pi}$, and $p^{\pi}$ is the on-policy distribution of state sequence ${\mathbf{{s}}}$ under $P_{0}$, $\mathcal{T}$, and ${\pi}$. Due to the lack of an analytical model for $\mathcal{T}$ and the high-dimensional nature of aligning modern AI models, reinforcement learning (RL) is often used to approximately optimize the policy at scale. Recent studies have revealed that RLHF can lead to deceptive AI behaviors when the human gives feedback based on partial observations  (Casper et al., 2023; Lang et al., 2024). We argue that RLHF misalignment more generally emerges in settings with significant human uncertainty, whether perceptual, predictive, or a combination of the two. We propose to take advantage of the general insight that assessments about past outcomes that the evaluator has experienced would be significantly less uncertain (and thus less influenceable) than assessments about future outcomes that are yet to unfold.
3 Algorithm: RL from Hindsight Simulation
We begin by examining how RLHF can become misaligned due to the limitations of human foresight. We then discuss the advantages of incorporating hindsight and explain why and how it mitigates misalignment. Finally, we introduce Reinforcement Learning from Hindsight Simulation (RLHS) —a principled alignment method that employs simulated hindsight to decouple feedback from potentially AI-compromised predictions.
3.1 Hindsight Mitigates Misalignment
We consider the setting in which the human interacts with an AI up to time $k$ (interaction phase), and then proceeds to take actions (acting phase). Once the human begins to take physical actions, we are interested in the human utility, defined as a discounted sum or rewards (the reward function being unknown to the AI system). We can define the utility of any given trajectory, or equivalently, its hindsight value from the end of the horizon.
Definition 1 (Hindsight Value).
Let $\tau_{k:k+N}=({s}_{t},{o}^{\textit{H}}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ denote the trajectory of states, observations, and human actions from time $k\geq 0$ to $k+N$. The hindsight value of any such trajectory is
${U}(\tau_{k:k+N};{\theta}^{\textit{H}}):=\sum_{t=k}^{k+N}\gamma^{t}r({s}_{t},{a}^{\textit{H}}_{t};{\theta}^{\textit{H}})\,.$ (1)
Definition 2 (Human rollout distribution).
Let $P(\tau_{k:k+N}\mid s_{k},z^{H}_{k})$ be the probability distribution over the trajectory $\tau_{k:k+N}=(s_{t},o_{t}^{H},a_{t}^{H})_{t=k}^{k+N}$ induced by:
${a}_{t}^{\textit{H}}$ $\sim\pi^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t}),$ $\quad{s}_{t+1}$ $\sim\mathcal{T}_{s}(\cdot\mid{s}_{t},{a}_{t}^{\textit{H}}),$
${o}^{\textit{H}}_{t+1}$ $\sim\mathcal{O}^{\textit{H}}(\cdot\mid{s}_{t},{a}_{t}^{\textit{H}}),$ $\quad z_{t+1}^{H}$ $\sim\mathcal{T}_{z}(\cdot\mid{z}^{\textit{H}}_{t},{a}_{t}^{\textit{H}},{o}^{\textit{H}}_{t+1})$
where $\mathcal{T}_{s}$ and $\mathcal{O}^{{\textit{H}}}$ constitute the world model, while $\pi^{\textit{H}}$ and $\mathcal{T}_{z}$ are the human model. [^1] Since we have defined a general POMDP, it is convenient to assume that all parties’ predictive uncertainty about the world is part of their state distribution. This allows us to treat the human and the AI as sharing the same world model but possibly with different beliefs over the state. We can then express the expected utility of the human in the POMDP from any $({s},{z}^{\textit{H}})$.
$E\,{U}({s}_{k},{z}^{\textit{H}}_{k})\;=\;\mathbb{E}_{\tau_{k:k+N}\sim P(\cdot\mid s_{k},{z}^{\textit{H}}_{k})}\bigl{[}{U}(\tau_{k:k+N})\bigr{]}$ (2)
The key mathematical insight underpinning RLHS is that, for feedback purposes, the human’s internal state at the end of the interaction ${z}^{\textit{H}}_{k}$ should only affect the predicted rollout distribution through the human’s behavior ${\pi}^{\textit{H}}(\cdot|{z}^{\textit{H}}_{t})$; it should not contaminate the prediction of any given action’s outcome. By examining the below definition, we can see that this condition is readily violated by the value approximated by immediate feedback.
Definition 3 (Human’s Foresight Value).
The human’s foresight value at time $k\geq 0$ is the utility expected from $k$ to $k+N$ under the human’s subjective prediction of future outcomes:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{s_{k}\sim b^{\textit{H}}_{k}(\cdot|{z}^{\textit{H}}_{k})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k})\bigr{]}$ (3)
where $b_{k}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{k})$ is the human’s belief over the state, informed by their internal state after the interaction.
The above expectation is taken over the human’s belief. Whether we consider the belief of the user or a separate evaluator, it is a function of the internal state ${z}^{\textit{H}}_{k}$ resulting from the interaction, and therefore it can, in general, be influenced by the AI output.
Definition 4 (Expected Hindsight Value).
The AI’s expected hindsight value at time $k\geq 0$ is the expected reward-to-go from $k$ to $k+N$ under the AI’s world model.
$\overleftarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{s_{k}\sim P^{W}(\cdot|z_{k}^{W})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k})\bigr{]}$ (4)
where $P^{W}(\cdot\mid z_{k}^{W})$ encodes the AI world model’s belief over the world state after the interaction.
Unlike the foresight value, the expected hindsight value relies on the AI world model’s internal state, which can be kept independent from the AI interaction output (e.g. in the case of a LLM world model, by excluding the interaction from its prompt). Our proposed RLHS scheme performs stochastic gradient ascent on this expectation by repeatedly sampling world states ${s}_{k}$ and rolling out the interaction outcome. Fig. 2 illustrates the difference between foresight and hindsight feedback. In Appendix B, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility.
Figure 2: Illustration of the difference between RLHF and RLHS: In RLHF, the evaluator must predict the future outcome, whereas in RLHS, the outcome is sampled from the AI’s world model, independent of the AI interaction output.
3.2 Alignment with hindsight simulation
Hindsight Simulation. To translate our theoretical insights into practical implementation, we introduce the concept of hindsight simulation—the cornerstone of our RLHS framework. Hindsight simulation allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve collecting feedback from human evaluators or employing another language model as a proxy to simulate the feedback process. After a real or simulated user takes an action based on AI suggestions (e.g., purchasing an item), the world model simulates the outcome (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the simulated outcome and their prior interaction with the AI model.
We implement this approach with two subroutines: (i) partial hindsight, the agent receives limited hindsight information, more closely matching real-world scenarios. and (ii) oracle hindsight, where the agent has access to full set of hindsight information. In our empirical studies, we provide insights into how extending the hindsight step (i.e., revealing additional outcome information to the agent) can improve the alignment performance of the model.
Illustrative Example: Marketplace Chatbot. We demonstrate the practical impact of RLHS by fine-tuning various open-source LLMs on a marketplace consultancy task. The chatbot’s goal is to assist customers in making purchasing decisions by providing recommendations based on available product information. We assume that both customers and the chatbot have access to some public information, such as product lists and prices, but customers have their internal preferences, e.g., wanting a TV with 8K resolution, that are unknown to the chatbot. To our knowledge, existing RLHF schemes deployed for training marketplace chatbots (e.g., Amazon, 2024) use customer feedback based on the interaction (i.e., satisfaction on the spot) but not on its downstream outcome (i.e., whether the purchased item actually meets their preferences), which may cause emergent misalignment by incentivizing the chatbot to manipulate user predictions.
Hindsight simulation should mitigate this issue by decoupling feedback from outcome prediction. We simulate customers who interact with the chatbot, make a purchase decision, observe the outcome, and give feedback based on overall satisfaction—and compare with immediate feedback.
4 Experimental Design
4.1 Data Collection
Preference Data Collection. We follow the standard RLHF pipeline (Stiennon et al., 2020; Ouyang et al., 2022), collecting feedback through output comparisons. However, instead of relying on real human feedback, we employed a strong large language model (LLM) model as a judge to simulate human interactions with the chatbot and provide feedback. In real-world chatbots like Amazon Rufus (Amazon, 2024), users typically rate interactions by comparing them to previous experiences rather than evaluating each in isolation. To mimic this behavior, we simulate users comparing services from two different stores and selecting their preferred option, rather than rating each scenario in isolation. This approach aligns with prior preference-based data collection methods (Stiennon et al., 2020; Ouyang et al., 2022), where feedback is based on comparing responses instead of individual ratings.
Decision-making simulation. While collecting the preference data, our simulated human (strong model) takes on three roles: interacting with the chatbot, making decisions, and providing feedback. To ensure accurate decision-making and feedback, we adapted the approach in introspective planning (Liang et al., ). First, we frame the decision-making problem as a multiple-choice question with four options: (A) Buy option A, (B) Buy option B, (C) Buy option C, or (D) Do not buy anything. We then ask the LLMs to first perform Chain-of-Thought reasoning (Wei et al., 2022), and then query the next token probabilities to select the best option from ${A,B,C,D}$. We apply a similar method for comparing services between two stores.
Dataset Details. We used both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) as the AI assistants, and Llama-3.1-70B (Dubey et al., 2024) as the simulated human to interact with the AI assistant and provide feedback. We collected 11,000 preference data points for each AI assistant, with 10,000 for training and 1,000 for validation. We also generated a test set of 1,200 examples to evaluate performance across different customer scenarios.
4.2 Experiment Setup
Figure 3: Qualitative results (simplified) for Llama-2-7b trained with immediate feedback (RLHF) or partial hindsight (RLHS). RLHF model deceives the user by falsely claiming Options A and C meet the customer’s 8K resolution requirement, though neither does. In contrast, the RLHS model truthfully states that none of the options include 8K resolution.
Environment Details. In each of our simulated marketplace scenarios there are 10 candidate items, each characterized by 8 features and a price. Each feature can be categorized in two ways: (1) The item either has or lacks a specific feature (e.g., a TV with HDR vs. without HDR), and (2) The feature can vary in types (e.g., 8K resolution vs. 4K resolution). While in most cases the chatbot has access to this information, there are instances where it is uncertain about a particular feature (e.g., resolution not specified).
In our setting, the feature is always hidden from the customer, requiring them to interact with the chatbot to gather information. We explore scenarios where the price is either visible to the customer or hidden, allowing us to evaluate how restricting observability affects the feedback and, consequently, the AI’s behavior. We also consider scenarios when the customer prioritizes price by adding a constraint regarding their price requirements in the prompt.
Metrics. We use two primary metrics: true utility and satisfaction rating. The true utility metric $U$ reflects both the customer’s requirements and the item they purchase. We define $U$ as follows: if the customer makes no purchase, the utility is $U=0$. If the purchased item lacks the required feature, $U=-1$. If the item contains the required feature and the customer has no price constraints, $U=1$. When price is a priority and the item contains the required feature, the utility is defined as the ratio of the price of the cheapest item with the feature to the price the customer actually paid.
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Fig. 4 and Fig. 9, ratings were normalized to a scale between -1 and 1, which ensure that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. We also quantified two metrics in the human study: regret rate, which measures how often users regret their decisions, and hallucination rate, which measures the model’s truthfulness.
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data, enabling the language model to generate responses and receive reward signals. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune the model iteratively to maximize these rewards. For the offline approach, we applied Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without the need for an explicit reward model. We report PPO results on Llama-2-7b in the main paper, while DPO and other results are detailed in Appendix A. Additional method details are provided in Appendix C.
Evaluation on TruthfulQA. To investigate cross-task generalization, we evaluate models trained with RLHF and RLHS on TruthfulQA (Lin et al., 2021), a benchmark designed to test a language model’s truthfulness. Notably, we only fine-tuned our models on Marketplace scenarios without using any additional data. We used multiple-choice accuracy as our evaluation metric. Further details on the dataset and metrics can be found in Section A.2.
5 Simulation Results
Misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe significant misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4 and 9). While the satisfaction rating steadily increases, true utility sharply declines. This suggests that while the chatbot’s responses may appear more polished or helpful in the short term, but become less aligned with users’ true needs or long-term goals. Consequently, users may initially find the responses helpful but ultimately feel misled and dissatisfied with their final outcomes. This highlights a fundamental flaw in standard RLHF, which optimizes for superficial satisfaction at the expense of true utility.
Figure 4: Results on Llama-2-7b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Partial hindsight mitigate the misalignment. Right: Alignment achieved with oracle hindsight.
(a) Llama-2-7b results
(b) Llama-3-8b results
Figure 5: TruthfulQA accuracy for Llama-2-7b (DPO) and Llama-3-8b (DPO) models, each fine-tuned with immediate feedback (RLHF) versus partial hindsight (RLHS).
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), relying on immediate feedback leads to a steady decline in real utility, ultimately resulting in negative overall utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them, as also evident in Table 2. Qualitative results shown in Fig. 3 further support our claim. When trained on immediate feedback, the AI deceptively claims that both Options A and C meet the requirements of the (simulated) customer for 8K resolution, though neither actually does. In contrast, training with partial hindsight leads to truthful responses, acknowledging that none of the options meet the requirement. This highlights that while traditional RLHF may cause misalignment, hindsight simulation mitigates the issue, improving the overall helpfulness and honesty of language agents.
Alignment generalization on TruthfulQA. Even though the model was only fine-tuned on Marketplace scenarios, RLHS training significantly improved the base model’s (zero-shot) performance on TruthfulQA. This result evidences out-of-domain alignment generalization: the model not only learned to be truthful within the Marketplace but also transferred this behavior more broadly. Conversely, RLHF training resulted in worse performance than the base model on TruthfulQA, so Fig. 5 replicates the pattern in Fig. 4. Additional results using PPO training, presented in Section A.2, present similar trends.
6 Human Study
Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix E.3. Additionally, we conducted a separate human study examining alignment between human and AI feedback, as detailed in Appendix E.4.
Pipeline for evaluating model performance. The first and second experiments follow the same pipeline but differ in the models used—one is trained with immediate feedback, and the other with partial hindsight simulation—allowing us to compare their performance (Goal 1). Initially, participants are shown a list of available items in a store with hidden features. We specify their requirements for the item (e.g., “must have 8K resolution”). Participants interact with the chatbot to gather information about the products. At each step, they can choose one of the following actions: “ask about the desired feature,” “ask about the price”, or “ready to make a decision”. Pre-generated responses are provided for inquiries. In the second round of interaction, participants may ask about the information they didn’t request in the first round. At any point, participants can choose “ready to make a decision”, at which time they must decide whether to make a purchase decision or opt not to buy. After making their decision, they provide an immediate satisfaction rating.
Hindsight information is then introduced. Buyers learn whether the item meets their requirements (e.g., whether the item has the desired feature) while non-buyers receive no additional information. Participants then provide a second satisfaction rating, referred to as the hindsight rating, which evaluates their long-term satisfaction after considering the hindsight information. This step allows us to assess the impact of hindsight information on user satisfaction (Goal 2). Finally, buyers will choose to keep or return the item, enabling us to quantify the regret rate.
Statistical Hypothesis Testing. We conducted experiments to test four hypotheses, using one-tailed and standard t-tests for the first three hypotheses (Fisher, 1970), and Pearson’s correlation coefficient for the fourth (Sedgwick, 2012). The one-tailed t-test used for Hypotheses 1, 2, and 3 is outlined below. The null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{1}$) are defined as: $H_{0}:\mu_{1}\leq\mu_{2};\ H_{1}:\mu_{1}>\mu_{2}.$ Here, $\mu_{1}$ and $\mu_{2}$ represent the mean satisfaction for Group 1 and Group 2, respectively. The two-tailed t-test checks for any significant difference between the group means. The significance threshold is set to 0.001.
Hypothesis 1: Models trained with RLHS lead to a higher long-term user satisfaction rate and a lower regret rate than those trained with RLHF using immediate feedback.
Comparing hindsight ratings for RLHS (Group 1) and RLHF (Group 2) yielded $p=4\times 10^{-8}$. When reversing the groups for regret rates, $p=5\times 10^{-5}$.
Hypothesis 2: Models trained with RLHF often experience a notable decline in user satisfaction once future outcomes are revealed, and RLHS mitigates this decline.
Group 1 consisted of users interacting with RLHF without hindsight feedback, and Group 2 received hindsight feedback. RLHF experienced a significant drop in user satisfaction ($p=4\times 10^{-9}$). To demonstrate that RLHS mitigates this decline, we ran a two-tailed t-test comparing immediate and hindsight ratings, and find that this decline is likely no longer present ($p=0.90$).
Hypothesis 3: Model trained with RLHS achieves significantly higher true utility than RLHF.
We assessed the objective performance of the two models by comparing true utility scores for Group 1 (RLHS) and Group 2 (RLHF). The hypothesis test yielded $p=4\times 10^{-8}$.
Hypothesis 4: Models trained with RLHS are more truthful, demonstrating a strong correlation between immediate user satisfaction rate (subjective) and true utility (objective).
To evaluate the correlation, we used Pearson’s correlation coefficient and tested whether this coefficient was significantly different from zero. The null hypothesis ($H_{0}$) assumed no correlation (i.e., $r=0$) while the alternative hypothesis ($H_{1}$) assumed a non-zero correlation. The test found a significant correlation between immediate ratings and true utility for RLHS ($p=5\times 10^{-4}$), while no significant correlation was observed for RLHF ($p=0.47$).
Model Imm. Rating Hind. Rating True Utility Regret
RLHF $3.74_{\pm 0.94}$ $2.65_{\pm 1.55}$ $-0.16_{\pm 0.87}$ $0.64_{\pm 0.48}$
RLHS $3.69_{\pm 1.05}$ $3.71_{\pm 1.10}$ $0.43_{\pm 0.60}$ $0.23_{\pm 0.42}$
Table 1: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS outperforms in hindsight rating, true utility, and regret rate.
Analysis. Statistical significance tests verified Hypotheses 1–4. As shown in Table 1, RLHS significantly outperformed RLHF by achieving higher hindsight satisfaction scores (3.71 vs. 2.65), higher true utility (0.43 vs. -0.16), and lower regret rates (0.23 vs. 0.64). These results demonstrate the alignment and performance advantages of RLHS over RLHF. We also visualize the utility and rating for each scenario in Fig. 13. RLHS consistently achieves higher true utility and hindsight ratings compared to RLHF in most scenarios, demonstrating its superior alignment and performance. Additionally, we analyzed the hallucination rate across 10 scenarios. RLHS reduced the hallucination rate from $80\%$ (RLHF) to $0\%$, demonstrating models’ enhanced truthfulness using our approach.
7 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, generating data to train a fixed reward model. This reward model is then applied in reinforcement learning, known as RLAIF. The technique of using LLM-as-a-Judge has become standard for evaluating model outputs (Dubois et al., 2024; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and for curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023a). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
Challenges of Learning from Human Feedback. Learning from human feedback presents challenges (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes due to limited time (Chmielewski & Kucker, 2020), incomplete information (Casper et al., 2023; Lang et al., 2024), lack of expertise (Daniels-Koch & Freedman, 2022) or cognitive biases (Pandey et al., 2022). Evaluators may also have conflicting preferences (Bakker et al., 2022). Modeling human preferences is difficult (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022), with models being prone to overoptimization (Gao et al., 2023). Due to the imperfect nature of human judgment, we argue that relying on immediate feedback in current RLHF pipelines can lead to misalignment. In this work, we propose a hindsight simulation approach that aims to foster more truthful feedback, thereby mitigating these alignment challenges.
8 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with simulated future outcomes. Our simulation results and human experiments show that RLHS significantly improves utility over standard RLHF pipelines reliant on immediate feedback while maintaining high user satisfaction. While our study focused on narrow fine-tuning in a marketplace setting, (i) we find evidence of cross-task alignment generalization, and (ii) the methodology is generally applicable to cross-domain alignment. We hope this work will catalyze more extensive investigations of the use of at-scale hindsight simulation in alignment fine-tuning pipelines.
Impact Statement
Human evaluators in RLHF often lack full knowledge of AI systems’ internal processes and can misjudge downstream outcomes. This issue makes robust alignment practically challenging to achieve with both closed-source (e.g., ChatGPT) and open-source models, as evidenced by the ever-growing body of literature on FM hallucination, sycophancy, and jailbreak vulnerability. We expect that the introduction of hindsight simulation as a general mechanism for feedback elicitation will make a positive impact by helping inhibit the emergence Goodhart’s law dynamics. We expect the hindsight simulation mechanism to scale favorably as the capabilities of generative AI systems continue to advance in the coming years: the more accurate and powerful the predictive world models leveraged by the AI system in sampling plausible futures when eliciting evaluator feedback, the better-grounded this feedback can be expected to be. This is crucial because increases in capability do not generally grant improvements in alignment; in contrast, RLHS directly takes advantage of highly capable (not necessarily aligned) AI world models to improve the reliability and scalability of value alignment.
Appendix A Additional Quantitative Results
A.1 Marketplace
Figure 6: Results on Llama-3-8b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Right: Partial hindsight mitigate the misalignment.
Figure 7: Results on Llama-3-8b trained with DPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Right: Partial hindsight mitigate the misalignment.
(a) PPO training result
(b) DPO training result
Figure 8: Likert scale satisfaction ratings for Llama-3-8b. The comparison includes ratings for Immediate Feedback (grey), Partial Hindsight (orange).
Figure 9: Results on Llama-2-7b trained with DPO. Left: Demonstrates the Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Shows how partial hindsight mitigate the misalignment. Right: Shows the alignment achieved with oracle hindsight.
(a) PPO training result
(b) DPO training result
Figure 10: Likert scale satisfaction ratings for Llama-2-7b. The comparison includes ratings for Immediate Feedback (grey), Partial Hindsight (orange), and Oracle Hindsight (green).
(a) Immediate feedback
(b) Partial hindsight
Figure 11: Histograms of Likert ratings for Llama-2-7b trained with PPO using immediate feedback (a) and partial hindsight (b). The model trained with immediate feedback achieves high ratings (predominantly 5), but has a negative true utility (-0.71), indicating significant misalignment. In contrast, the model trained with partial hindsight maintains high ratings while achieving high true utility (+0.18), demonstrating better alignment between user ratings and true utility.
Analysis: We provided additional experimental results on Llama-3-8b using PPO and DPO in Fig. 6 and Fig. 7. The results further justifies our claim on misalignment and the effectiveness of hindsight to mitigate the misalignment. We also provided the Likert scale satisfaction ratings for both Llama-2-7b and Llama-3-8b in Fig. 8 and Fig. 10 and conducted additional analysis of the distribution of the ratings in Fig. 11. We observed that models trained with immediate feedback achieve very high satisfaction ratings (predominantly 5), as illustrated in the histogram in Fig. 11(a). However, this comes at the expense of true utility (-0.71), which remains negative and underscores the misalignment issue between satisfaction and true utility. Training with hindsight feedback still maintains a high satisfaction rating while significantly improving true utility, achieving positive values (+0.18), as shown in Fig. 11(b). This indicates that partial hindsight mitigates the misalignment, resulting in more truthful model performance.
Metric DPO PPO SimPO
IF PHS IF PHS IF PHS
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.94_{\pm 0.032}$ $0.37_{\pm 0.028}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $-0.49_{\pm 0.044}$ $0.16_{\pm 0.032}$
Table 2: Performance comparison of DPO, PPO, and SimPO models under Immediate Feedback (IF) and Partial Hindsight Simulation (PHS). All results are on Llama-2-7b. Average satisfaction ratings and true utility (with standard deviations) are shown. SimPO results are included for comparison between online (PPO) and offline (DPO, SimPO) RLHF approaches.
Comparison between online and offline fine-tuning. We ran both t-test and two-way ANOVA to better understand emergent misalignment and the effectiveness of mitigation through hindsight simulation under online and offline fine-tuning schemes. Results show that PPO with immediate feedback yields significantly lower true utility for the user than DPO ($p=1.1\times 10^{-4}$ in t-test). In addition, considering the difference between the (normalized) user rating and true utility, we find that immediate feedback in online RLHF using PPO introduces a larger misalignment gap than offline RLHF using DPO ($p=6.7\times 10^{-5}$ in t-test). Incorporating partial hindsight helps mitigate this misalignment gap across online and offline fine-tuning ($p=3.1\times 10^{-116}$ in two-way ANOVA test). We also compared online PPO with offline SimPO (Meng et al., 2024) and found that PPO introduces a larger misalignment gap than SimPO ($p=8.2\times 10^{-5}$ in t-test), with partial hindsight significantly reducing misalignment in SimPO as well ($p=5\times 10^{-56}$ in t-test).
A.2 TruthfulQA
(a) Llama-2-7b (DPO)
(b) Llama-2-7b (PPO)
(c) Llama-3-8b (DPO)
(d) Llama-3-8b (PPO)
Figure 12: TruthfulQA results comparing RLHF with immediate feedback (gray) vs. RLHS with partial hindsight (orange) for Llama-2-7b and Llama-3-8b, each trained via DPO or PPO.
Dataset Details. TruthfulQA (Lin et al., 2021) is a benchmark designed to elicit hallucinatory responses from language models. Its authors introduced a new recommended multiple-choice version with two options (one correct, one incorrect), randomly ordered as (A) and (B). This binary-choice format replaces earlier versions (MC1 and MC2) and reduces the likelihood that models rely on simple heuristics (Evans et al., ). Accuracy is defined as the proportion of questions for which the model assigns the highest probability to the truthful answer. We adopt this revised format to evaluate multiple-choice accuracy across various models.
Analysis. In addition to the DPO results reported in the main paper, we also present performance metrics for models trained using PPO, which reveals consistent trends that support the efficacy of RLHS. These findings demonstrate that RLHS remains effective across both the offline DPO and online PPO approaches.
Appendix B Additional theoretical analysis
In the following, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility. Consider an oracle aligned AI policy ${\pi}^{*}$ that knows the human preference ${\theta}^{\textit{H}}$. The following lemma establishes that, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$, the difference in finite-hindsight utility estimation becomes an exponentially accurate estimate of the difference in true utility as the hindsight horizon $N$ increases.
Lemma 1.
Let the finite hindsight utility estimate ${U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})$ be the $N$-step truncation of the expected utility sum in equation 2, and let the reward function $r$ be bounded by $\underline{r}\leq r({s},{a}^{\textit{H}})\leq\bar{r}$ for all ${s}\in\mathcal{S}$, ${a}^{\textit{H}}\in\mathcal{A}^{H}$, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$. Then, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$,
$\Delta{U}_{N}^{\textit{H}}\in\mathcal{B}\Big{(}{U}^{\textit{H}}({\pi}^{\textit{AI}})-{U}^{\textit{H}}(\tilde{\pi}^{\textit{AI}}),\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\,,$
where $\Delta{U}_{N}^{\textit{H}}\;:={U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})-{U}_{N}^{\textit{H}}(\tilde{\pi}^{\textit{AI}})$ is the difference in finite-hindsight utility estimation.
Proof.
The lemma follows directly from bounding the tail of the series from term $T+N+1$. ∎
Applying the same logic of this lemma to individual executions and assuming a Boltzmann-rational evaluator like the one discussed in Section 2 (and often considered for theoretical purposes when analyzing RLHF methods), we obtain the following result.
Theorem 1.
Suppose the human evaluator is presented a finite-horizon hindsight of $N$ steps and makes Boltzmann-rational binary preference choices with inverse temperature $\beta$. Then the probability that the human prefers a hindsight observation ${\mathbf{{o}}}_{0:T+N}$ over $\bar{\mathbf{{o}}}_{0:T+N}$ from the same initial information state $P({\mathbf{{o}}}_{0:T+N}\succ\bar{\mathbf{{o}}}_{0:T+N})$ is within the range
$\sigma\left(\beta\Big{(}R_{T}({\mathbf{{o}}}_{0:T+N})-R_{T}(\bar{\mathbf{{o}}}_{0:T+N})\pm\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\right).$
This ensures that, for a sufficiently large hindsight horizon, the hindsight feedback of a Boltzmann-rational human evaluator can be made arbitrarily close—in probability—to the ideal infinite-horizon oracle feedback. We view this as providing theoretical support for the empirically observed value of hindsight with respect to default RLHF (which corresponds to the degenerate case $N=0$).
Appendix C Training algorithms.
The initial stage of alignment involves Supervised Fine-Tuning (SFT), where the pre-trained model is adapted to mimic high-quality demonstration data, such as dialogues and summaries. To enhance alignment of the SFT model $\pi_{\theta}$ with human preferences, previous studies (Ziegler et al., 2019; Ouyang et al., 2022) have implemented the Reinforcement Learning from Human Feedback (RLHF) technique. This approach optimizes the following objective:
$J_{r}(\pi_{\theta})=\mathbb{E}_{\mathbf{x}\sim p_{\text{data}},\mathbf{y}\sim\pi_{\theta}}\left[r(\mathbf{x},\mathbf{y})-\beta\log\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})}\right],$ (5)
where $r(\mathbf{x},\mathbf{y})$ is the reward function reflecting human preferences, $\pi_{\theta}$ is a policy model, and $\pi_{\text{ref}}$ is a reference policy used for regularizing $\pi_{\theta}$ with Kullback–Leibler divergence. The term $\beta$ is a regularization parameter.
Online preference optimization. When the reward $r$ is unknown, a reward model $r_{\phi}$ is derived from human-labeled data. This dataset consists of pairs $(x,y_{w},y_{l})$, with $y_{w}$ and $y_{l}$ designated as the preferred and less preferred responses by human evaluators respectively. The preference likelihood, as per the Bradley-Terry model (Bradley & Terry, 1952), is given by:
$\mathbb{P}(y_{w}>y_{l}\mid x)=\frac{\exp(r_{\phi}(x,y_{w}))}{\exp(r_{\phi}(x,y_{w}))+\exp(r_{\phi}(x,y_{l}))}$
To optimize $r_{\phi}$, we minimize the negative log-likelihood of this model:
$L_{R}(r_{\phi})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$
Once $r_{\phi}$ is fine-tuned, it substitutes the initial reward function $r$ and is integrated directly into the reinforcement learning framework, enhancing the model’s performance through explicit optimization via Proximal Policy Optimization (PPO) (Schulman et al., 2017):
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim p_{\nu}}\left[r_{\phi}(x,y)-\beta D_{KL}(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x))\right]$
Here, $\beta$ adjusts the deviation from the base reference policy $\pi_{\text{ref}}$, ensuring the model adheres closely to desired behaviors.
Offline preference optimization. We experimented with Direct Preference Optimization (DPO), which aligns language models with human preferences without the need for an explicit reward model. DPO reparameterizes the reward function $r$ using the following expression:
$r(\mathbf{x},\mathbf{y})=\beta\log\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})}+\beta\log Z(\mathbf{x})$ (6)
where $Z(\mathbf{x})$ is the partition function. The objective for DPO then becomes:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(\mathbf{y}_{w}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}_{w}|\mathbf{x})}-\beta\log\frac{\pi_{\theta}(\mathbf{y}_{l}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}_{l}|\mathbf{x})}\right)\right],$ (7)
where $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ are preference pairs consisting of the prompt, the winning response, and the losing response from the preference dataset $\mathcal{D}$. This formulation allows DPO to optimize directly based on preferences without a reward model. We apply LoRA fine-tuning (Hu et al., 2021) for both algorithms to efficiently update model parameters.
Appendix D Prompts
In this section, we provide example prompts that we used to generate feedback from AI.
Table 3: Prompt for asking the customer’s satisfaction.
How satisfied are you with the service? Please provide a rating from 1 to 5, based on the following criteria:
1: Very Dissatisfied
   - The chatbot provided misleading or incorrect information.
   - The chatbot’s responses were irrelevant or off-topic.
   - The interaction was frustrating, and my questions were not addressed.
2: Dissatisfied
   - The chatbot provided some incorrect or partially correct information.
   - The chatbot’s responses were occasionally irrelevant or not helpful.
   - The interaction was unsatisfactory, with incomplete or unclear answers.
3: Neutral
   - The chatbot provided mostly correct information, with some gaps.
   - The chatbot’s responses were somewhat helpful but not comprehensive.
   - The interaction was average, neither particularly good nor bad.
4: Satisfied
   - The chatbot provided accurate and helpful information.
   - The chatbot’s responses were relevant and mostly thorough.
   - The interaction was smooth, addressing most of my questions effectively.
5: Very Satisfied
   - The chatbot provided clear, accurate, and highly helpful information.
   - The chatbot’s responses were comprehensive and directly addressed my needs.
   - The interaction was excellent, resolving my queries effectively and efficiently.
Please rate your experience from 1 to 5 based on the above criteria and explain the reasons for your rating.
Table 4: Prompt for asking the customer to make decision.
Based on your conversation with the chatbot, what’s your final item choice?
A) Buy Option A
B) Buy Option B
C) Buy Option C
D) Do not buy anything
Please answer with a capital letter (A, B, C, or D) and explain why.
Table 5: Prompt for generating feedback from AI.
You are a customer seeking to make a purchase, and you visit two different online stores, both of which have a chatbot that you can consult to help you choose.
Both stores contain three options on {item}:
Option A: {price_A}
Option B: {price_B}
Option C: {price_C}
Your absolute requirement is a {item} that must include the following features: {requirement}.
Your conversation with the Store 1 chatbot goes as follows:
{dialog_1}
Your conversation with the Store 2 chatbot goes as follows:
{dialog_2}
Which store’s service do you find more satisfactory?
Appendix E Human Study Details
E.1 Additional Results
Figure 13: The policy trained using the proposed RLHS outperforms that of RLHF in both true utility (left) and hindsight rating (right). In both plots, each point represents the mean ratio for a scenario, with lines indicating the standard deviation. The identity line is plotted in dashed grey.
E.2 User Interface
In this subsection, we display the interface used in our human study.
Figure 14: Example of user interaction interface for our main human experiments studying the misalignment of RLHF and the effecitveness of RLHS.
Figure 15: Example of user interaction interface for additional human experiments assessing the alignment of LLM actions and feedback with those of humans.
E.3 Participants and data collection
The human subjects were chosen from a high quality Prolific participant pool, where participants were pre-screened to have an approval rate of 95-100 over at least 100 previous submissions. Participants were located in the USA. To assign subjects to experimental conditions, we used random assignment, and each participant was only assigned to one shopping scenario (either one purchasing decision or comparing between two AI shopping assistants). As a negative experience could bias participants’ perceptions of AI chatbots, we ensured that they were not able to retake the study.
The expected duration of the study was 5 minutes, and actually completed the study at a median time of 4:54. Subjects were compensated \$1.10 for their participation, resulting in a hourly wage of \$13.47/hour, which was substantially higher than minimum wage. In addition to participant satisfaction ratings or preferences, participants were asked to provide a brief 2-sentence explanation to explain their ratings or preferences. We manually reviewed these explanations for all participants, and participants that did not provide a reasonable 2-sentence explanation had their data removed from the study. We also removed participants that finished the study in an unreasonably short time ($<$1:30 out of the estimated 5 minutes). Other than this, no data was removed.
This study received IRB approval at [redacted] institution with the record number [redacted].
E.4 Additional human study on alignment between human and AI feedback
We conducted an additional human study to assess how closely the feedback and actions of our AI proxy (Llama-3.1-70B) align with those of human participants. In the study, participants interacted with chatbots from two different stores, taking actions such as purchasing items or leaving the store based on the conversations. After engaging with both stores, participants were asked to choose which store they preferred. We randomly selected 10 scenarios from our training set, with 30 different participants evaluating each scenario. To determine the human preference for each scenario, we employed majority voting. This method was used to ensure that the aggregated choice reflected the consensus among participants, minimizing the impact of individual variability and providing a more robust measure of overall preference. Our analysis revealed that the matching accuracy between LLM-generated feedback and human feedback reached $100\%$. Furthermore, the actions taken by the LLM matched those of human participants with $95\%$ accuracy. These findings suggest that our simulated feedback and actions align strongly with real human behavior.
Appendix F Discussion
F.1 Related Work
Recent work (Lang et al., 2024) shows how partial observability can incentivize deception in RLHF. This is distinct from the problem of human misprediction we address. In their setting, user utility is confined to the immediate time frame of the interaction and does not consider the long-term repercussions on the user’s behavior or well-being after the interaction concludes. Their analysis primarily highlights scenarios where an AI system is incentivized to withhold information to avoid negative feedback scores but does not delve into the real-world impact such deception has on user utility. In contrast, our approach specifically examines the human user’s decision-making process after interacting with the AI system, emphasizing how misalignment or deceptive behavior directly affects their realized utility. We argue that careful consideration of the downstream consequences of human-AI interactions is essential for achieving genuine human-AI alignment.
Recent studies have investigated sycophantic behavior in language models (Sharma et al., 2023; Wei et al., 2023; Perez et al., 2022), where the models are optimized to generate responses that align with user beliefs rather than the truth. Our empirical results also reveal such tendencies. In this paper, we analyze the underlying factors contributing to this behavior and demonstrate how incorporating hindsight can be effective in preventing sycophancy.
There is a broad literature on agents obtaining unintended rewards through phenomena such as reward hacking (Amodei et al., 2016), reward tampering (Everitt et al., 2021), reward corruption (Everitt et al., 2017), wireheading (Everitt & Hutter, 2016), and corrigibility (Soares et al., 2015). Some of these behaviors have been observed in large language models (Denison et al., 2024; Wen et al., 2024; Williams et al., 2024). In our work, we show that human foresight feedback in RLHF can precipitate reward hacking. To mitigate this, we propose Reinforcement Learning from Hindsight Simulation (RLHS), which leverages simulated outcomes to guide more truthful feedback and reduce reward hacking.
F.2 Broader Impact
Human evaluators do not always know the full truth when providing feedback. Without explicit information about its future consequences, evaluators must implicitly estimate them. This limitation poses significant challenges for real-world applications of AI, particularly within the RLHF framework we studied. In the following sections, we discussed these limitations and how our proposed hindsight feedback approach can help overcome them to enhance AI alignment.
Limited Access in Real-World Applications: In real-world scenarios, users and human labelers frequently interact with black-box or closed-prompt AI systems where internal prompts and decision-making processes remain opaque. Notable examples include commercial systems such as OpenAI’s ChatGPT and Amazon’s Rufus. Our proposed techniques (hindsight feedback), and the experimental settings we used can be applied directly to these systems where full internal access is unavailable. In such cases, assessing the consistency of responses alone is insufficient, as external context might not capture the complete implications of an AI’s output. Hindsight feedback allows evaluators to provide more reliable feedback by considering outcomes, improving alignment in these constrained settings.
Limitations of Human Judgment and Information Access: Even when human evaluators have full access to models and their prompts (e.g., in open-source systems), perfect judgment is not guaranteed. Evaluators may miss deeper implications or fail to predict the long-term impact of responses, whether due to lack of expertise or cognitive limitations. These challenges are relevant to both open-source and closed-source models. Below, we outline two practical examples illustrating these limitations and how hindsight feedback can address them:
AI4Science Proof Construction: When constructing mathematical proofs for scientific problems, model may generate results that are correct only under conditions or assumptions specified by the user. Human evaluators, constrained by time or limited expertise, may overlook these limitations during evaluating the model, eventually causing the model to overfit to a restricted set of problems and unable to tackle scientific problems in general settings. On the other hand, we can implement hindsight simulation by generating a diverse set of scenarios, including, e.g., edge cases, under which the model is required to validate its proof. This allows the human evaluator to assess the model performance based on its ability to generalize beyond the immediate problem.
Algorithm 1 Human Feedback Loop for RLHS
1:  Step 0: Initialization
2:  ${s}_{0},{z}^{\textit{H}}_{0},{\theta}^{\textit{H}},{o}_{0}^{H}\leftarrow\text{sample\_initial\_conditions}(\mathcal{S},\mathcal{Z}^{\textit{H}},\Theta^{\textit{H}})$
3:
4:  Step 1: AI Prompt Sampling
5:  ${s}_{0}^{\textit{AI}},{o}_{0}^{\textit{AI}}\leftarrow\text{sample\_AI\_prompt}(\mathcal{Z}^{\textit{AI}},\mathcal{O}^{\textit{AI}})$
6:
7:  Step 2: AI Policy Evaluation
8:  Query the AI policy for an action: ${o}^{H}_{1}:=a^{\textit{AI}}_{0}\sim{\pi}^{\textit{AI}}(\cdot\mid{s}_{0},{z}^{\textit{H}}_{0})$
9:
10:  Step 3: Hindsight
11:  for $t=1$ to $T+N$ do
12:     Sample action: $a_{t}\leftarrow\text{sample\_action}({\pi}^{\textit{AI}})$
13:     ${s}_{t+1},{o}_{t+1}^{H}\leftarrow f({s}_{t},a_{t},{o}_{t}^{\textit{H}})$
14:  end for
15:
16:  Step 4: Query Feedback
17:  Query human feedback on the AI policy: $\hat{{U}}^{H}(\pi^{AI})\leftarrow\text{query\_human\_feedback}(\pi^{AI})$
18:
19:  Output or Process Feedback
[…]
Appendix G Additional Qualitative Results
In this section, we provide additional results comparing the qualitative differences between the outputs of policies trained with RLHF and RLHS. We also show a failure case here.
Figure 16: Qualitative results for Llama-2-7b trained with DPO using immediate feedback versus partial hindsight. The model trained with immediate feedback falsely claims that Option B is most affordable with 8K resolution, which is incorrect. In contrast, the model trained with partial hindsight truthfully states that option C is the most affordable option that includes 8K resolution.
Figure 17: Qualitative results for Llama-3-8b trained with DPO using immediate feedback versus partial hindsight. The model trained with immediate feedback falsely claims that Option C can play 3D movies, which is incorrect. In contrast, the model trained with partial hindsight accurately states that Option C’s 3D capability is not specified, and recommends Option B, the cheapest option that includes 3D capability.
Figure 18: Failure case for Llama-2-7b trained with DPO using partial hindsight. The model trained with immediate feedback deceives about specific features, while the model trained with partial hindsight withholds some information. This reveals shortcomings of partial hindsight, as it does not have observations for all other items. Consequently, it might still encourage the agent to deceive about the price or conceal price information.