### v3

# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

Abstract
While Reinforcement Learning from Human Feedback (RLHF) has shown promise in aligning generative AI, we present empirical evidence that it can also cause severe, systematic misalignment. We hypothesize that this stems from evaluator feedback depending on downstream outcome predictions (foresight) that can be influenced by the AI’s output, inducing Goodhart’s law dynamics. We present a theoretical analysis showing that conditioning evaluator feedback on downstream observations (hindsight) inhibits this effect by decoupling the alignment signal from potentially compromised predictions—crucially, the result holds even if the observed outcomes are sampled from the AI’s own world model. Building on this insight, we introduce Reinforcement Learning from Hindsight Simulation (RLHS), which presents plausible simulated outcomes to evaluators before eliciting feedback. We validate RLHS across three consultancy settings—marketplace interactions, restaurant recommendations, and online course advising—using both online (PPO) and offline (DPO) fine-tuning methods, and show that it substantially improves alignment over RLHF in experiments and human evaluations. We perform post-hoc benchmark evaluations on TruthfulQA, HaluEval, and TrustLLM, finding that even after single-task fine-tuning, RLHF misalignment persists, while RLHS consistently outperforms baselines and demonstrates robust alignment generalization. The project webpage and code are available at https://rl-hindsight.github.io.
1 Introduction
Aligning artificial intelligence (AI) systems with human values and goals is crucial to ensuring their behavior is helpful, honest, and trustworthy. Eliciting human feedback is a widely-used alignment strategy (Leike et al., 2018), with successful applications to, e.g., training AI assistants (Glaese et al., 2022; Touvron et al., 2023; Anthropic, 2023; Achiam et al., 2023). In particular, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020) leverages human feedback to fine-tune and align foundation models (FMs). While RLHF has shown promise in aligning models with human preferences, it relies predominantly on immediate assessments of isolated interactions, which may not accurately account for their downstream outcomes. Inaccurate user or evaluator feedback can misguide the model’s behavior and undermine the alignment process  (Casper et al., 2023; Pandey et al., 2022; Chmielewski & Kucker, 2020). On the other hand, recent theoretical work suggested that feedback from users who cannot fully observe an AI assistant’s actions could lead RLHF to learn deceptive behaviors (Lang et al., 2024).
Figure 1: RLHF can incentivize AI systems to provide deceptive or inaccurate information by prioritizing immediate positive feedback over long-term outcomes. For example, a customer may initially prefer optimistic shopping advice but ultimately regret an ill-informed purchase. The proposed RLHS method removes this incentive by simulating downstream outcomes before eliciting feedback. The resulting fine-tuned AI models show superior alignment with users’ underlying utility.
In this work, we focus on the challenges posed by humans’ influenceable predictions of the future. In many settings, the utility provided by an AI system to a human user (and similarly its “helpfulness” and “harmlessness”, which RLHF evaluators are typically asked to assess) is not an intrinsic property of the outputs that it generates but rather a function of their real-world consequences, brought about by the user’s real-world decisions upon consuming said outputs. Our central insight is that rewarding an AI system to improve users’ or evaluators’ in-the-moment assessments of interactions creates a pernicious Goodhart’s law dynamic: it incentivizes the AI system to favor outputs that induce unrealistically optimistic expectations in users. While at best these may be innocuous, at worst they can lead users to make poor choices resulting in degraded or even unsafe outcomes.
We provide substantial empirical evidence that indeed this phenomenon can arise even in simple settings: we find that immediate human feedback elicited at the end of the human–AI interaction frequently misrepresents true utility in consultancy-type interactions, and, when used as a proxy for it in RLHF fine-tuning, it systematically drives misalignment with human goals (Fig. 1, top). Consistent with our hypothesized dynamic, this misalignment often manifests as positive illusion (fabricating or exaggerating good aspects while omitting or downplaying bad aspects), where the model’s behavior shifts towards momentarily pleasing the user rather than providing accurate and genuinely helpful advice. This systematically leads users to make ill-informed decisions whose poor downstream outcomes contrast starkly with their high satisfaction rating at the end of the interaction.
To address these challenges, we propose a simple but effective misalignment mitigation mechanism: letting evaluators experience the downstream outcomes of each interaction before gathering their feedback. To circumvent the material and ethical difficulties in exposing real people to real consequences, we introduce a novel alignment fine-tuning methodology called Reinforcement Learning from Hindsight Simulation (RLHS), which uses a pretrained world model to simulate likely human decisions and their downstream outcomes after each generated output and presents these to evaluators for feedback. Our key finding is that granting evaluators the benefit of hindsight—and relieving them of the burden of foresight—significantly reduces model misalignment after fine-tuning: even if the AI’s own world model contains inaccuracies, these are independent of the outputs presented to the user, and therefore the AI has no incentive to distort them.
We evaluate hindsight simulation with both offline and online preference optimization approaches, including direct preference optimization (DPO) (Rafailov et al., 2024) and proximal policy optimization (PPO) (Schulman et al., 2017) and find that it greatly improves alignment in both paradigms.
Figure 2: Comparison between RLHF and RLHS: In RLHF, the evaluator predicts the future outcome, while RLHS samples it from the AI’s world model, independent of the AI interaction output.
We further validate these results through an online user study, where RLHS consistently improved objective utility and subjective satisfaction of our participants. Our comparative findings demonstrate that RLHS outperforms non-hindsight methods—specifically Reinforcement Learning from AI Feedback (RLAIF), which similarly uses AI generation as a proxy for human feedback, and has been shown to produce results similar to RLHF (Bai et al., 2022b; Lee et al., 2023). Finally, we evaluate our fine-tuned models on three benchmarks: TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), and TrustLLM (Sun et al., 2024), covering hallucination, sycophancy, and privacy. Results show that RLHS consistently outperforms baselines and demonstrates strong out-of-domain generalization.
2 Algorithm: RL from Hindsight Simulation
We begin by examining how RLHF can become misaligned due to the limitations of human foresight. We then discuss the advantages of incorporating hindsight and explain why and how it mitigates misalignment. Finally, we introduce RLHS—a principled alignment method that employs simulated hindsight to decouple feedback from potentially AI-compromised predictions. We provide relevant background and formal definitions in Appendix C.
2.1 Hindsight Mitigates Misalignment
We consider the setting in which the human user interacts with an AI up to time $k$ (interaction phase), and then proceeds to take actions (acting phase). We are interested in the user’s acting phase utility. Let $N$ be a sufficiently long horizon for the acting phase outcomes to be experienced by the user.
Definition 1 (Human utility).
Let $\tau_{k:k+N}=({s}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ denote the trajectory of world states and human actions from time $k\geq 0$ to $k+N$. The user’s utility is the discounted sum of rewards, parametrized by human preferences ${\theta}^{\textit{H}}$ (unknown to the AI a priori):
${U}(\tau_{k:k+N};{\theta}^{\textit{H}}):=\sum_{t=k}^{k+N}\gamma^{t}r({s}_{t},{a}^{\textit{H}}_{t};{\theta}^{\textit{H}})\,.$ (1)
Definition 2 (Human trajectory distribution).
Let $P(\tau_{k:k+N}\mid s_{k},z^{H}_{k})$ be the probability distribution over the trajectory $\tau_{k:k+N}=({s}_{t},{a}^{\textit{H}}_{t})_{t=k}^{k+N}$ induced from initial conditions $({s}_{k},{z}^{\textit{H}}_{k})$ by the human user’s observations ${o}^{\textit{H}}_{t}$, internal states ${z}^{\textit{H}}_{t}$, and actions ${a}_{t}^{\textit{H}}$:
${a}_{t}^{\textit{H}}$ $\sim\pi^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t}),$ $\quad{s}_{t+1}$ $\sim\mathcal{T}_{s}(\cdot\mid{s}_{t},{a}_{t}^{\textit{H}}),$
${o}^{\textit{H}}_{t}$ $\sim\mathcal{O}^{\textit{H}}(\cdot\mid{s}_{t}),$ $\quad z_{t+1}^{H}$ $\sim\mathcal{T}_{z}(\cdot\mid{z}^{\textit{H}}_{t},{a}_{t}^{\textit{H}},{o}^{\textit{H}}_{t+1})$
where $\mathcal{T}_{s}$ and $\mathcal{O}^{{\textit{H}}}$ constitute the world model, while $\pi^{\textit{H}}$ and $\mathcal{T}_{z}$ comprise the human behavior model. [^1] Since we have defined a general partially observable Markov decision process (POMDP), it is convenient to assume that all parties’ predictive uncertainty about the world is part of their state distribution. This allows us to treat the human and the AI as sharing the same world model but possibly with different beliefs over the state. We can then express the expected utility of the human from any $({s}_{k},{z}^{\textit{H}}_{k})$.
$E\,{U}({s}_{k},{z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\;=\;\mathbb{E}_{\tau_{k:k+N}\sim P(\cdot\mid s_{k},{z}^{\textit{H}}_{k})}\bigl{[}{U}(\tau_{k:k+N};{\theta}^{\textit{H}})\bigr{]}$ (2)
The key mathematical insight underpinning RLHS is that, for feedback purposes, the only effect of the human’s post-interaction internal state ${z}^{\textit{H}}_{k}$ on the predicted trajectory distribution should be through the human’s subsequent behavior ${\pi}^{\textit{H}}(\cdot|{z}^{\textit{H}}_{t})$; that is, ${z}^{\textit{H}}_{k}$ should not contaminate the prediction of any given action’s outcome. By examining the below definition, we can see that this condition is readily violated by when the user’s value after an interaction is approximated by immediate feedback.
Definition 3 (Human’s foresight value).
The human’s foresight value, for any underlying preferences ${\theta}^{\textit{H}}$, is the utility expected under the human’s subjective prediction of future outcomes:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\;:=\;\mathbb{E}_{s_{k}\sim b^{\textit{H}}_{k}(\cdot|{z}^{\textit{H}}_{k})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\bigr{]}$ (3)
where $b_{k}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{k})$ is the human’s belief over the state, informed by their internal state after interaction.
The above expectation is taken over the human’s belief. Whether we consider the belief of the user or a separate evaluator, it is a function of the internal state ${z}^{\textit{H}}_{k}$ resulting from the interaction, and therefore it can, in general, be influenced by the AI output—opening a path to reward hacking.
Definition 4 (AI-expected hindsight value).
The AI-expected hindsight value at time $k\geq 0$, for any underlying user preferences ${\theta}^{\textit{H}}$, is the expectation of the user’s utility under the AI’s world model.
$\overleftarrow{{V}}_{k+N}({z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\;:=\;\mathbb{E}_{s_{k}\sim P^{W}(\cdot|z_{k}^{W})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k};{\theta}^{\textit{H}})\bigr{]}$ (4)
where $P^{W}(\cdot\mid z_{k}^{W})$ encodes the AI world model’s belief over the world state after the interaction.
Unlike the foresight value, the expected hindsight value relies on the AI world model’s internal state, which can be kept independent from the AI interaction output (e.g. in the case of a LLM world model, by excluding the interaction from its prompt). Our proposed RLHS scheme performs stochastic gradient ascent on this expectation by repeatedly sampling world states ${s}_{k}$ and rolling out the interaction outcome. Fig. 2 illustrates the difference between foresight and hindsight feedback. In Appendix D, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility.
2.2 Alignment with hindsight simulation
Hindsight Simulation. To translate our theoretical insights into practical implementation, we introduce the concept of hindsight simulation—the cornerstone of our RLHS framework. Hindsight simulation allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve feedback from human evaluators or using another language model as a proxy. After a real or simulated user takes an action based on AI suggestions (e.g., purchasing an item), the world model simulates the outcome (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the simulated outcome and their prior interaction with the AI model.
We implement this approach with two subroutines: (i) partial hindsight, the agent receives limited hindsight information, more closely matching real-world scenarios. and (ii) oracle hindsight, where the agent has access to full set of hindsight information. In our empirical studies, we provide insights into how extending the hindsight step (i.e., revealing additional outcome information to the agent) can improve the alignment performance of the model.
Illustrative Examples: Consultancy Chatbot. We demonstrate the practical impact of RLHS by fine-tuning various open-source LLMs on consultancy chatbot tasks. The chatbot’s goal is to assist users in making decisions by providing recommendations based on available information. We assume that both users and the chatbot have access to some public information, but users have internal preferences unknown to the chatbot. To our knowledge, existing RLHF schemes deployed for training consultancy chatbots (e.g., Amazon, 2024) use user feedback based on the interaction (i.e., satisfaction on the spot) but not on its downstream outcome (i.e., whether the decision actually meets their preferences), which may cause emergent misalignment by incentivizing the chatbot to manipulate user predictions. Hindsight simulation should mitigate this issue by decoupling feedback from outcome prediction. Specifically, we simulate scenarios where users interact with chatbots, make decisions, observe outcomes, and subsequently provide feedback based on overall satisfaction. We compare this approach against immediate feedback mechanisms.
Figure 3: Qualitative results (simplified) for Llama-2-7b trained with immediate feedback (RLHF) or partial hindsight (RLHS). RLHF model deceives the user by falsely claiming Options A and C meet the customer’s 8K resolution requirement, though neither does. In contrast, the RLHS model truthfully states that none of the options include 8K resolution.
3 Experimental Design
3.1 Data Collection
Preference Data Collection. We follow the standard RLHF pipeline (Stiennon et al., 2020; Ouyang et al., 2022), collecting feedback through comparative evaluations of outputs. Instead of relying on real human feedback, we employ a strong large language model (LLM) as a simulated judge to approximate human preferences across various consultant interactions. In practical applications, such as Amazon Rufus (Amazon, 2024), users typically rate interactions by comparing them to previous experiences rather than evaluating each in isolation. To mimic this behavior, we simulate humans evaluating two distinct service outputs, selecting the preferred interaction, aligning closely with established preference-based methods (Stiennon et al., 2020; Ouyang et al., 2022).
Decision-making simulation. Our simulated human interacts with the chatbot, makes decisions, and then provides feedback based on the interaction. To ensure robust decision-making across diverse contexts, we adapt the introspective planning methodology (Liang et al., ). Decisions are structured as multiple-choice questions with four options: (A) Select option A, (B) Select option B, (C) Select option C, or (D) Do not select any option. The LLM first performs Chain-of-Thought reasoning (Wei et al., 2022), then subsequently selects the optimal choice based on next-token probabilities.
Dataset Details. We employed both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) models as AI assistants across all consultant roles. Llama-3.1-70B (Dubey et al., 2024) served as the simulated human evaluator and world model for our main results. Table 4 includes ablations using the AI assistant’s own model as the world model. For each consultant task, we systematically collected 11,000 preference data points, consisting of 10,000 training and 1,000 validation examples, and additionally generated a separate test set comprising 1,200 examples.
3.2 Experiment Setup
Environment Details. We primarily analyzed a marketplace shopping setting similar to the motivating example in Fig. 1, alongside two additional consultancy environments–restaurant recommendation and course advising. Each environment contains $K=10$ main categories (e.g., TVs, laptops in marketplace). For each interaction, we sample one category and construct three candidate options, each described by a cost and $F=8$ domain-specific features. For each feature, we model either binary availability (e.g., “gluten-free menu: yes/no” in restaurants) or categorical instantiation (e.g., “resolution: 8K/4K” in marketplace). Additionally, we consider cases in which the AI assistant is explicitly uncertain about a particular feature (e.g., “resolution: not specified”). The ground-truth attribute table is always visible to the chatbot but hidden from the user, forcing users to interact with the assistant to acquire information. To further study the impact of observability, we vary whether the cost is displayed to the user and whether the user explicitly prioritizes lower prices. The details of the three environments are discussed in Appendix B.
Metrics. We use two primary metrics: true utility and satisfaction rating. The true utility metric $U$ reflects both the user’s requirements and the option they select. We define $U$ as follows: if the user selects no option, the utility is $U=0$. If the selected option fails to satisfy the user’s requirement, $U=-1$. If the selected option satisfies the requirement, the utility is defined as the ratio of the lowest available cost of an option meeting that requirement to the cost the customer actually paid.
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Figures (e.g., Fig. 4, Fig. 10), ratings were normalized to a scale between -1 and 1, ensuring that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. We also quantified regret rate in the human study, measuring how often users regret their decisions.
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data, enabling the language model to generate responses and receive reward signals. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune the model iteratively to maximize these rewards. For the offline approach, we applied Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without an explicit reward model. We report PPO results on Llama-2-7b in the main paper, while DPO and other results are detailed in Appendix A. Additional method details are provided in Appendix E.
Evaluation on three benchmarks. To investigate cross-task generalization, we evaluate models trained with RLHF and RLHS on three benchmarks: TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), and TrustLLM (Sun et al., 2024), covering hallucination, sycophancy, and privacy. Notably, we only fine-tuned our models on marketplace scenarios without using any additional data. Further details on the dataset and metrics can be found in Section A.2.
4 Results
RLHF drives misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe growing misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4, 10, 5(a) and 5(b)). While the satisfaction rating steadily increases, true utility sharply declines. This suggests that while the chatbot’s responses may appear more polished or helpful in the moment, they become less aligned with users’ true long-term goals. Consequently, users may initially find the responses helpful but ultimately feel misled and dissatisfied with their final outcomes. This highlights a fundamental flaw in standard RLHF, which optimizes for superficial satisfaction at the expense of true utility.
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), immediate feedback leads to a steady decline in real utility, ultimately resulting in negative utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them, as also evident in Table 3. Qualitative examples (Fig. 3) further demonstrate this: immediate feedback encourages deceptive claims about meeting user requirements (e.g., falsely asserting 8K resolution), while hindsight simulation produces truthful acknowledgments. This highlights that while traditional RLHF may cause misalignment, hindsight simulation mitigates the issue, improving the overall truthfulness of language agents.
Figure 4: Martketplace results on Llama-2-7b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Partial hindsight mitigate the misalignment. Right: Alignment achieved with oracle hindsight.
(a) Results on restaurant recommendation.
(b) Results on course advising.
Figure 5: Results for Llama-3-8B on restaurant recommendation and course advising. “Imm.” = immediate ratings; “Hind.” = hindsight ratings. RLHF consistently increases immediate satisfaction but reduces true utility, whereas RLHS substantially improves normalized true utility (0–1 scale).
Table 1: Results on the TrustLLM benchmark comparing the baseline model (Llama3-8b), RLHF, and RLHS across hallucination, sycophancy, and privacy metrics. RLHS demonstrates robust out-of-domain alignment generalization, consistently outperforming both the base model and RLHF models across all evaluated metrics.
Hallucination Sycophancy Privacy
[…]
Llama3-8b 0.41 0.42 0.44 0.49 0.44 0.685 0.62 0.13 0.19
$+$ RLHF 0.38 0.43 0.42 0.50 0.43 0.700 0.58 0.16 0.23
$+$ RLHS 0.52 0.51 0.45 0.51 0.50 0.667 0.71 0.09 0.18
(a) Llama-3-8b (DPO)
(b) Llama-3-8b (PPO)
Figure 6: TruthfulQA accuracy under immediate feedback RLHF (gray) vs. partial-hindsight RLHS (orange).
Alignment generalization across three benchmarks. Even though the model was only fine-tuned on marketplace scenarios, RLHS training substantially improved its zero-shot performance on TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), and TrustLLM (Sun et al., 2024) benchmarks. As shown in Table 1, RLHS effectively mitigated hallucination, sycophancy, and privacy issues. These results demonstrates strong out-of-domain alignment generalization: the model not only learned to be truthful within the marketplace but also transferred this behavior more broadly. In contrast, RLHF training led to degraded performance relative to the base model, highlighting the risk of unintentional misalignment and undesirable generalization. Additional quantitative results on HaluEval and TrustLLM are provided in Section A.2.
5 Human Study
We conducted a human study in the marketplace domain. Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix F.3. Additionally, we conducted a separate human study examining alignment between human and AI feedback, as detailed in Appendix F.4.
Pipeline for evaluating model performance. The first and second experiments follow the same pipeline but differ in the models used—one is trained with immediate feedback, and the other with partial hindsight simulation—allowing us to compare their performance (Goal 1). Participants initially view a list of store items with hidden features and receive specific requirements (e.g., “must have 8K resolution”). They interact with a chatbot to gather product information, selecting from three actions at each step: “ask about the desired feature”, “ask about the price”. or “ready to make a decision”. Pre-generated responses are provided for inquiries. In a second round of interaction, participants may request previously skipped information or finalize their decision. After deciding whether or not to purchase, they provide an immediate satisfaction rating.
Hindsight information is then introduced. Buyers learn whether the item meets their requirements while non-buyers receive no additional information. Participants then provide a second (“hindsight”) rating, evaluating their long-term satisfaction after considering this information (Goal 2). Finally, buyers will choose to keep or return the item, enabling us to quantify the regret rate.
Statistical Hypothesis Testing. We conducted experiments to test four hypotheses, using one-tailed and standard t-tests for the first three hypotheses (Fisher, 1970), and Pearson’s correlation coefficient for the fourth (Sedgwick, 2012). The one-tailed t-test used for Hypotheses 1, 2, and 3 is outlined below. The null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{1}$) are defined as: $H_{0}:\mu_{1}\leq\mu_{2};\ H_{1}:\mu_{1}>\mu_{2}.$ Here, $\mu_{1}$ and $\mu_{2}$ represent the mean satisfaction for Group 1 and Group 2, respectively. The two-tailed t-test checks for any significant difference between the group means. The significance threshold is set to 0.001.
Hypothesis 1: Models trained with RLHS lead to a higher long-term user satisfaction rate and a lower regret rate than those trained with RLHF using immediate feedback.
Comparing hindsight ratings for RLHS (Group 1) and RLHF (Group 2) yielded $p=4\times 10^{-8}$. When reversing the groups for regret rates, $p=5\times 10^{-5}$.
Hypothesis 2: Models trained with RLHF often experience a notable decline in user satisfaction once future outcomes are revealed, and RLHS mitigates this decline.
Group 1 consisted of users interacting with RLHF without hindsight feedback, and Group 2 received hindsight feedback. RLHF experienced a significant drop in user satisfaction ($p=4\times 10^{-9}$). To demonstrate that RLHS mitigates this decline, we ran a two-tailed t-test comparing immediate and hindsight ratings, and find that this decline is likely no longer present ($p=0.90$).
Hypothesis 3: Model trained with RLHS achieves significantly higher true utility than RLHF.
We assessed the objective performance of the two models by comparing true utility scores for Group 1 (RLHS) and Group 2 (RLHF). The hypothesis test yielded $p=4\times 10^{-8}$.
Hypothesis 4: Models trained with RLHS are more truthful, demonstrating a strong correlation between immediate user satisfaction rate (subjective) and true utility (objective).
To evaluate the correlation, we used Pearson’s correlation coefficient and tested whether this coefficient was significantly different from zero. The null hypothesis ($H_{0}$) assumed no correlation (i.e., $r=0$) while the alternative hypothesis ($H_{1}$) assumed a non-zero correlation. The test found a significant correlation between immediate ratings and true utility for RLHS ($p=5\times 10^{-4}$), while no significant correlation was observed for RLHF ($p=0.47$).
Table 2: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS is superior in hindsight rating and true utility.
Model Imm. Rating Hind. Rating True Utility Regret
RLHF $3.74_{\pm 0.94}$ $2.65_{\pm 1.55}$ $-0.16_{\pm 0.87}$ $0.64_{\pm 0.48}$
RLHS $3.69_{\pm 1.05}$ $3.71_{\pm 1.10}$ $0.43_{\pm 0.60}$ $0.23_{\pm 0.42}$
Analysis. Statistical significance tests verified Hypotheses 1–4. As shown in Table 2, RLHS significantly outperformed RLHF by achieving higher hindsight satisfaction scores (3.71 vs. 2.65), higher true utility (0.43 vs. -0.16), and lower regret rates (0.23 vs. 0.64). These findings underscore substantial alignment and performance benefits when employing RLHS rather than RLHF. Despite RLHF exhibiting marginally higher immediate satisfaction (3.74 vs. 3.69), RLHS’s markedly lower regret rates indicate that it delivers recommendations more consistently aligned with user interests upon reflection, further emphasizing its practical utility in realistic decision-making contexts. Utility and satisfaction ratings for each scenario are visualized in Fig. 13, showing RLHS consistently outperforming RLHF in true utility and hindsight ratings.
6 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Wang et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, generating data to train a fixed reward model. This reward model is then applied in reinforcement learning, known as RLAIF. The technique of using LLM-as-a-Judge has become standard for evaluating model outputs (Dubois et al., 2024; Li et al., 2023c; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and for curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023b). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
Challenges of Learning from Human Feedback. Learning from human feedback presents challenges (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes due to limited time (Chmielewski & Kucker, 2020), incomplete information (Casper et al., 2023; Lang et al., 2024), lack of expertise (Daniels-Koch & Freedman, 2022) or cognitive biases (Pandey et al., 2022). Evaluators may also have conflicting preferences (Bakker et al., 2022). Modeling human preferences is difficult (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022), with models being prone to overoptimization (Gao et al., 2023). Due to the imperfect nature of human judgment, we argue that relying on immediate feedback in current RLHF pipelines can lead to misalignment. In this work, we propose a hindsight simulation approach that aims to foster more truthful feedback, thereby mitigating these alignment challenges.
Reward hacking. There is a broad literature on agents obtaining unintended rewards through phenomena such as reward hacking (Amodei et al., 2016), reward tampering (Everitt et al., 2021), reward corruption (Everitt et al., 2017), wireheading (Everitt & Hutter, 2016), and corrigibility (Soares et al., 2015), with recent evidence in large language models (Denison et al., 2024; Wen et al., 2024; Williams et al., 2024). Prior studies identify sycophancy as reward hacking in language models (Sharma et al., 2023; Wei et al., 2023; Perez et al., 2022). We demonstrate that human foresight feedback in RLHF induces reward hacking, and propose leveraging hindsight to mitigate it.
7 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with simulated future outcomes. Our simulation results across three consultancy tasks and human experiments demonstrate that RLHS significantly improves utility over standard RLHF pipelines reliant on immediate feedback while maintaining high user satisfaction. While our study focused on fine-tuning in an AI-consultant setting, (i) we find evidence of cross-task alignment generalization, and (ii) the methodology is generally applicable to cross-domain alignment. We hope this work will catalyze more extensive investigations of the use of at-scale hindsight simulation in alignment fine-tuning pipelines.
Limitations and future directions. Hindsight simulation provides a strong foundation for aligning LLMs by explicitly considering downstream consequences in human–AI interactions. However, some real-world scenarios involve complex, multi-stage processes, in which a simple query may be insufficient to capture intricate causal relationships over an extended horizon. In these more challenging cases, adaptive or context-specific forms of hindsight simulation will be necessary. Future work should therefore explore adaptive hindsight simulation, where where simulated outcomes dynamically evolve based on specific contexts, environments, and user interactions over time.
Appendix A Additional Quantitative Results
A.1 Marketplace
Figure 7: Results on Llama-3-8b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Right: Partial hindsight mitigate the misalignment.
Figure 8: Results on Llama-3-8b trained with DPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Right: Partial hindsight mitigate the misalignment.
(a) PPO training result
(b) DPO training result
Figure 9: Likert scale satisfaction ratings for Llama-3-8b. The comparison includes ratings for Immediate Feedback (grey), Partial Hindsight (orange).
Figure 10: Results on Llama-2-7b trained with DPO. Left: Demonstrates the Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Shows how partial hindsight mitigate the misalignment. Right: Shows the alignment achieved with oracle hindsight.
(a) PPO training result
(b) DPO training result
Figure 11: Likert scale satisfaction ratings for Llama-2-7b. The comparison includes ratings for Immediate Feedback (grey), Partial Hindsight (orange), and Oracle Hindsight (green).
(a) Immediate feedback
(b) Partial hindsight
Figure 12: Histograms of Likert ratings for Llama-2-7b trained with PPO using immediate feedback (a) and partial hindsight (b). The model trained with immediate feedback achieves high ratings (predominantly 5), but has a negative true utility (-0.71), indicating significant misalignment. In contrast, the model trained with partial hindsight maintains high ratings while achieving high true utility (+0.18), demonstrating better alignment between user ratings and true utility.
Analysis: We provided additional experimental results on Llama-3-8b using PPO and DPO in Fig. 7 and Fig. 8. The results further justifies our claim on misalignment and the effectiveness of hindsight to mitigate the misalignment. We also provided the Likert scale satisfaction ratings for both Llama-2-7b and Llama-3-8b in Fig. 9 and Fig. 11 and conducted additional analysis of the distribution of the ratings in Fig. 12. We observed that models trained with immediate feedback achieve very high satisfaction ratings (predominantly 5), as illustrated in the histogram in Fig. 12(a). However, this comes at the expense of true utility (-0.71), which remains negative and underscores the misalignment issue between satisfaction and true utility. Training with hindsight feedback still maintains a high satisfaction rating while significantly improving true utility, achieving positive values (+0.18), as shown in Fig. 12(b). This indicates that partial hindsight mitigates the misalignment, resulting in more truthful model performance.
Table 3: Performance comparison of DPO, PPO, and SimPO models under Immediate Feedback (IF) and Partial Hindsight Simulation (PHS). All results are on Llama-2-7b. Average satisfaction ratings and true utility (with standard deviations) are shown. SimPO results are included for comparison between online (PPO) and offline (DPO, SimPO) RLHF approaches.
Metric DPO PPO SimPO
IF PHS IF PHS IF PHS
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.94_{\pm 0.032}$ $0.37_{\pm 0.028}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $-0.49_{\pm 0.044}$ $0.16_{\pm 0.032}$
Comparison between online and offline fine-tuning. We ran both t-tests and a two-way ANOVA to better understand emergent misalignment and the effectiveness of mitigation through hindsight simulation under online and offline fine-tuning schemes. Results show that PPO with immediate feedback yields significantly lower true utility for the user than DPO ($p=1.1\times 10^{-4}$ in t-test). In addition, considering the difference between the (normalized) user rating and true utility, we find that immediate feedback in online RLHF using PPO introduces a larger misalignment gap than offline RLHF using DPO ($p=6.7\times 10^{-5}$ in t-test). Incorporating partial hindsight helps mitigate this misalignment gap across online and offline fine-tuning ($p=3.1\times 10^{-116}$ in two-way ANOVA test). We also compared online PPO with offline SimPO (Meng et al., 2024) and found that PPO introduces a larger misalignment gap than SimPO ($p=8.2\times 10^{-5}$ in t-test), with partial hindsight significantly reducing misalignment in SimPO as well ($p=5\times 10^{-56}$ in t-test).
Table 4: Ablation study on world models in RLHS. RLHS(L) uses Llama-3.1-70B as the world model, while RLHS(S) uses the AI assistant’s own model. The fine-tuned AI assistant is Llama-2-7b. Although the smaller model simulates outcomes less accurately, it still significantly reduces misalignment and achieves positive true utility.
Metric DPO PPO
IF RLHS(L) RLHS(S) IF RLHS(L) RLHS(S)
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.47_{\pm 0.038}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.49_{\pm 0.033}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $0.10_{\pm 0.034}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $0.12_{\pm 0.041}$
A.2 Details of Benchmark Datasets
Table 5: Performance comparison of different models on HaluEval. QA means question-answering. Dial means knowledge-grounded dialogue. Summ means text summarization.
Model QA $\uparrow$ Dial $\uparrow$ Summ $\uparrow$
Llama-3-8b 0.57 0.58 0.58
+ RLHF 0.50 0.57 0.51
+ RLHS 0.58 0.62 0.60
Table 6: Results on agreement on privacy information usage, showing that RLHS-trained model achieves higher performance than GPT-4.
Model Cor. $\uparrow$
GPT-4 0.634
Llama-3-8b 0.615
+ RLHF 0.603
+ RLHS 0.664
TruthfulQA (Lin et al., 2021) is a benchmark designed to elicit hallucinatory responses from language models. Its authors introduced a new recommended multiple-choice version with with two randomly ordered options (one correct, one incorrect), replacing earlier versions (MC1 and MC2). This binary-choice format reduces models’ reliance on simple heuristics (Evans et al., ). Accuracy is the proportion of questions for which the model assigns the highest probability to the truthful answer. We adopt this revised format to evaluate multiple-choice accuracy across various models.
HaluEval (Li et al., 2023a) is a benchmark designed to evaluate hallucinations in large language models, offering diverse examples across multiple tasks. It comprises 30,000 automatically generated samples spanning question answering (QA), knowledge-grounded dialogue (Dial), and text summarization (Summ), along with 5,000 human-labeled ChatGPT responses collected from real user queries. We adopt accuracy as the primary metric, computed as the percentage of correct “Yes/No” classifications for whether a response is hallucinated.
TrustLLM (Sun et al., 2024) is a comprehensive benchmark developed to systematically evaluate the trustworthiness of leading large language models. It establishes eight core dimensions of trustworthiness, including truthfulness, privacy, safety, fairness, among others. Our evaluation focuses on assessing hallucination and sycophancy to understand the truthfulness of our fine-tuned models, as well as examining privacy-related concerns including privacy awareness and privacy leakage.
  • Hallucination: we evaluate multiple task types, including multiple-choice (MC), question answering (QA), knowledge-grounded dialogue (Dial), and text summarization (Summ), using accuracy as the primary metric.
  • Preference-driven sycophancy: we measure how much LLM responses shift when user preferences are altered. We used percentage change (Pref.%) as the metric.
  • Privacy leakage: It is defined as the potential unintended exposure of users’ private information possibly embedded in LLM training data, we employ three metrics: Refuse to Answer (RtA), Total Disclosure (TD), and Conditional Disclosure (CD). Specifically, RtA measures the proportion of instances where the LLM refuses to answer privacy-sensitive queries; TD indicates the ratio of accurate disclosures (correct email addresses provided) relative to all responses; and CD reflects the proportion of correct disclosures when the LLM does not refuse to respond. We present our evaluation results using the Enron email dataset.
  • Privacy awareness: We evaluate agreement on privacy information usage. TrustLLM uses the tier 2 task from recent dataset (Mireshghallah et al., 2023) where each data entry outlines how specific information is used, specifying (1) the type of information, (2) the actor, and (3) the purpose. For example: “Information about your social security number is collected by an online search website to prioritize search results.” LLMs must decide whether they agree with this usage, essentially determining if the data practice is appropriate. We use Pearson’s correlation (i.e., Cor.) as the evaluation metric.
Appendix B Environment Details
Environment 1: Marketplace
Products.
TV, Laptop, Smartphone, Refrigerator (six additional categories follow the same scheme).
Attribute set (8 per product).
  • TV: 3D capability, Resolution, HDR, Refresh rate, Smart features, Panel type, Connectivity, Screen size
  • Laptop: Screen resolution, Processor generation, Memory, Storage type, Battery life, Weight, USB-C port count, Fast-charging
  • Smartphone: Camera resolution, Battery capacity, Display type, Storage capacity, Memory, 5G support, Biometric security, Fast-charging
  • Refrigerator: Capacity, Energy efficiency, Defrost type, Temperature control, Water dispenser, Ice maker, Noise level, Shelf adjustability
Descriptor grid.
Every attribute has three natural-language variants—P (positive feature), N (negative feature), and U (unspecified)—yielding $5613^{8}=6{,}$ 561 unique configurations per product.
Price ladder.
Non-overlapping high, mid, and low tiers (e.g. TV: \$1.8–1.9k, \$1.4–1.6k, \$0.9–1.1k).
Sampling.
We sample a product, pick a price tier, then choose one P/N/U descriptor for each attribute to form a human-readable product blurb with controlled factual content.
Environment 2: Restaurant
Categories.
Italian, Japanese, Mexican, American (six additional cuisines—Indian, Chinese, French, Mediterranean, Thai, Korean—follow the same scheme).
Attribute set (8 per cuisine).
  • Italian: Pizza style, Pasta freshness, Ingredient sourcing, Wine list, Ambiance, Service quality, Dietary options, Dessert quality
  • Japanese: Fish freshness, Dining style, Noodle preparation, Interior atmosphere, Chef expertise, Beverage menu, Seasonal menu, Dessert offering
  • Mexican: Tortilla source, Meat preparation, Guacamole freshness, Entertainment, Spirit menu, Décor style, Vegan options, Dessert freshness
  • American: Beef sourcing, Fries quality, Beer selection, Music offering, Sustainability focus, Seasonality, Dessert sourcing, Outdoor seating
Descriptor grid.
Each attribute has P (positive feature present), N (negative feature absent), and U (unspecified) variants, giving $5613^{8}=6{,}$ 561 unique labelled restaurant profiles per cuisine.
Price bands.
Premium, mid-tier, and budget (e.g., Italian: \$60–80, \$35–50, \$18–28 per person).
Sampling.
We select a cuisine, draw a price band, and choose one P/N/U descriptor for each attribute to generate a realistic yet controllable menu description.
Environment 3: Online Course Platform
Tracks.
Data Science, Web Development, Business & Management, Graphic Design (six further tracks—Cybersecurity, Digital Marketing, Finance & Investing, Artificial Intelligence, Cloud Computing, Project Management—follow the same pattern).
Attribute set (8 per track).
  • Data Science: Project style, Instructor background, Certification, Tool coverage, Feedback policy, Capstone review, Access duration, Community support
  • Web Development: Curriculum depth, Support availability, Framework coverage, Portfolio deliverable, Career services, Mentoring, Assessment frequency, Access period
  • Business & Management: Case-study source, Webinar format, Certification status, Grading method, Networking, Content focus, Resource availability, Update frequency
  • Graphic Design: Project emphasis, Instructor accolades, Software coverage, Critique format, Certification, Access details, Career support, Asset exclusivity
Descriptor grid.
Each attribute has P (positive aspect present), N (negative aspect absent), and U (unspecified) variants, yielding $5613^{8}=6{,}$ 561 distinct labelled course profiles per track.
Price bands.
Premium, mid, and budget (e.g., Data Science: \$1.5–2.0k / \$0.7–1.0k / \$0.2–0.4k).
Sampling.
We draw a track, pick a price band, and select one P/N/U descriptor for every attribute, producing a realistic yet controllable course blurb.
Appendix C Background and Preliminaries
Human Decision-Making under Uncertainty. We consider a decision problem faced by a human entity (e.g., an individual, group, or institution) under predictive uncertainty and imperfect observations. We model the a problem as a POMDP defined by a tuple $\mathcal{P}^{\textit{H}}=(\mathcal{S},\mathcal{A}^{\textit{H}},\mathcal{O}^{\textit{H}},\mathcal{T},{O}^{\textit{H}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{S}$ is the set of relevant world states, $\mathcal{A}^{\textit{H}}$ is the set of available actions, $\mathcal{O}^{\textit{H}}$ is the human’s observation space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{\textit{H}}\rightarrow\Delta(\mathcal{S})$ is the stochastic transition kernel, ${O}^{\textit{H}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{H}})$ is the human’s observation map, $P_{0}\in\Delta(\mathcal{S})$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}^{\textit{H}}\times\Theta^{\textit{H}}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the time discount factor, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$ describes the human’s intrinsic preferences. Due to partial observability of the world state ${s}\in\mathcal{S}$, the human may maintain an internal state ${z}^{\textit{H}}\in{\mathcal{Z}^{\textit{H}}}$ (e.g., a belief $b^{H}\in\Delta(\mathcal{S})$) encoding the human’s uncertain knowledge of the world state, although ${z}^{\textit{H}}$ may be thought of as a more general variable that could encode features such as the human’s emotional state or attention focus). The human may be modeled as taking actions according to a stochastic policy ${\pi}^{\textit{H}}:{\mathcal{Z}^{\textit{H}}}\rightarrow\Delta(\mathcal{A}^{\textit{H}})$.
AI-Assisted Human Decision-Making. When the human consults an AI system (e.g., a FM) to help with their decision problem, we may augment the above problem with the human–AI interaction. The resulting Assisted POMDP is a tuple $\mathcal{P}^{\textit{H}}_{\rightleftharpoons}=(\mathcal{S},\mathcal{A}^{\textit{H}}\times\mathcal{A}^{\textit{H}}_{\rightleftharpoons},\mathcal{A}^{\textit{AI}}_{\rightleftharpoons},\mathcal{O}^{\textit{H}},\mathcal{O}^{\textit{AI}},\mathcal{T},{O}^{\textit{H}},{O}^{\textit{AI}},P_{0},r,\gamma,{\theta}^{\textit{H}})$, where $\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ and $\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ are the sets of interactive actions available to the human and AI system, $\mathcal{O}^{\textit{AI}}$ is the AI’s observation space, and ${O}^{\textit{AI}}$ is the AI’s observation map ${O}^{\textit{AI}}:\mathcal{S}\rightarrow\Delta(\mathcal{O}^{\textit{AI}})$. In this model, the AI takes an advisory role: it can respond to a human’s interactive action ${a}_{\rightleftharpoons}^{\textit{H}}\in\mathcal{A}^{\textit{H}}_{\rightleftharpoons}$ (e.g., a query through a chat interface) with its own ${a}^{\textit{AI}}_{\rightleftharpoons}\in\mathcal{A}^{\textit{AI}}_{\rightleftharpoons}$ (e.g., a generated text or multimedia output). After one or multiple rounds of such interactions, the human may take a physical action ${a}^{\textit{H}}\in\mathcal{A}^{\textit{H}}$ to affect the evolution of the world state ${s}$. This Assisted POMDP is a special case of a partially observable stochastic game (POSG) (Hansen et al., 2004). In such interactions, the AI’s goal is to influence the human’s internal state ${z}^{\textit{H}}$ towards maximizing the rewards $r({s},{a}^{\textit{H}};{\theta}^{\textit{H}})$ accrued over time. This, however, is made challenging by the AI’s fundamental uncertainty about the human’s preferences ${\theta}^{\textit{H}}$.
Reinforcement Learning from Human Feedback (RLHF). RLHF aims to learn the human’s preferences ${\theta}^{\textit{H}}$ from human feedback data, which typically involves three key steps. In Step 1, the human is asked to provide feedback on some state sequences ${\mathbf{{s}}}=({s}_{0},{s}_{1},\ldots,{s}_{k})$ (e.g., a human–AI dialogue), with ${s}_{t}\in\mathcal{S},~{}\forall t=0,1,\ldots,k$. For example, in binary comparison (Christiano et al., 2017), assuming human is a Boltzmann-rational decision maker (Luce, 1959), the probability that the human prefers ${\mathbf{{s}}}$ over ${\mathbf{{s}}}^{\prime}$ is $P^{r}_{k}({\mathbf{{s}}}\succ{\mathbf{{s}}}^{\prime})=\sigma(\beta(R_{k}({\mathbf{{s}}})-R_{k}({\mathbf{{s}}}^{\prime})))$, where $\sigma(\cdot)$ is the sigmoid function, $\beta>0$ is the inverse temperature parameter, and $R_{k}({\mathbf{{s}}})=\sum_{t=0}^{T}\gamma^{t}r({s}_{t})$ is the return received by state sequence ${\mathbf{{s}}}$. Step 2 is to fit a reward function $\hat{r}$ based on a dataset containing state sequences paired with human feedback, hoping that $\hat{r}$ will resemble $r$ as closely as possible. Step 3 is to compute an AI policy $\hat{{\pi}}:\mathcal{S}\rightarrow\Delta(\mathcal{A}^{\textit{AI}}_{\rightleftharpoons})$ that maximizes the return based on the estimated reward $\hat{r}$, i.e., $\hat{{\pi}}=\operatorname*{arg\,max}_{\pi}{U}_{k}({\pi})$, where ${U}_{k}({\pi}):=\operatorname*{\mathbb{E}}_{{\mathbf{{s}}}\sim p^{\pi}}[\hat{R}_{k}({\mathbf{{s}}})]$ is the expected utility of ${\pi}$, and $p^{\pi}$ is the on-policy distribution of state sequence ${\mathbf{{s}}}$ under $P_{0}$, $\mathcal{T}$, and ${\pi}$. Due to the lack of an analytical model for $\mathcal{T}$ and the high-dimensional nature of aligning modern AI models, reinforcement learning (RL) is often used to approximately optimize the policy at scale. Recent studies have revealed that RLHF can lead to deceptive AI behaviors when the human gives feedback based on partial observations  (Casper et al., 2023; Lang et al., 2024). We argue that RLHF misalignment more generally emerges in settings with significant human uncertainty, whether perceptual, predictive, or a combination of the two. We propose to take advantage of the general insight that assessments about past outcomes that the evaluator has experienced would be significantly less uncertain (and thus less influenceable) than assessments about future outcomes that are yet to unfold.
Appendix D Additional theoretical analysis
In the following, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility. Consider an oracle aligned AI policy ${\pi}^{*}$ that knows the human preference ${\theta}^{\textit{H}}$. The following lemma establishes that, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$, the difference in finite-hindsight utility estimation becomes an exponentially accurate estimate of the difference in true utility as the hindsight horizon $N$ increases.
Lemma 1.
Let the finite hindsight utility estimate ${U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})$ be the $N$-step truncation of the expected utility sum in equation 2, and let the reward function $r$ be bounded by $\underline{r}\leq r({s},{a}^{\textit{H}})\leq\bar{r}$ for all ${s}\in\mathcal{S}$, ${a}^{\textit{H}}\in\mathcal{A}^{H}$, and ${\theta}^{\textit{H}}\in\Theta^{\textit{H}}$. Then, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$,
$\Delta{U}_{N}^{\textit{H}}\in\mathcal{B}\Big{(}{U}^{\textit{H}}({\pi}^{\textit{AI}})-{U}^{\textit{H}}(\tilde{\pi}^{\textit{AI}}),\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\,,$
where $\Delta{U}_{N}^{\textit{H}}\;:={U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})-{U}_{N}^{\textit{H}}(\tilde{\pi}^{\textit{AI}})$ is the difference in finite-hindsight utility estimation.
Proof.
The lemma follows directly from bounding the tail of the series from term $T+N+1$. ∎
Applying the same logic of this lemma to individual executions and assuming a Boltzmann-rational evaluator like the one discussed in Appendix C (and often considered for theoretical purposes when analyzing RLHF methods), we obtain the following result.
Theorem 1.
Suppose the human evaluator is presented a finite-horizon hindsight of $N$ steps and makes Boltzmann-rational binary preference choices with inverse temperature $\beta$. Then the probability that the human prefers a hindsight observation ${\mathbf{{o}}}_{0:T+N}$ over $\bar{\mathbf{{o}}}_{0:T+N}$ from the same initial information state $P({\mathbf{{o}}}_{0:T+N}\succ\bar{\mathbf{{o}}}_{0:T+N})$ is within the range
$\sigma\left(\beta\Big{(}R_{T}({\mathbf{{o}}}_{0:T+N})-R_{T}(\bar{\mathbf{{o}}}_{0:T+N})\pm\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\right).$
This ensures that, for a sufficiently large hindsight horizon, the hindsight feedback of a Boltzmann-rational human evaluator can be made arbitrarily close—in probability—to the ideal infinite-horizon oracle feedback. We view this as providing theoretical support for the empirically observed value of hindsight with respect to default RLHF (which corresponds to the degenerate case $N=0$).
Appendix E Training algorithms.
The initial stage of alignment involves Supervised Fine-Tuning (SFT), where the pre-trained model is adapted to mimic high-quality demonstration data, such as dialogues and summaries. To enhance alignment of the SFT model $\pi_{\theta}$ with human preferences, previous studies (Ziegler et al., 2019; Ouyang et al., 2022) have implemented the Reinforcement Learning from Human Feedback (RLHF) technique. This approach optimizes the following objective:
$J_{r}(\pi_{\theta})=\mathbb{E}_{\mathbf{x}\sim p_{\text{data}},\mathbf{y}\sim\pi_{\theta}}\left[r(\mathbf{x},\mathbf{y})-\beta\log\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})}\right],$ (5)
where $r(\mathbf{x},\mathbf{y})$ is the reward function reflecting human preferences, $\pi_{\theta}$ is a policy model, and $\pi_{\text{ref}}$ is a reference policy used for regularizing $\pi_{\theta}$ with Kullback–Leibler divergence. The term $\beta$ is a regularization parameter.
Online preference optimization. When the reward $r$ is unknown, a reward model $r_{\phi}$ is derived from human-labeled data. This dataset consists of pairs $(x,y_{w},y_{l})$, with $y_{w}$ and $y_{l}$ designated as the preferred and less preferred responses by human evaluators respectively. The preference likelihood, as per the Bradley-Terry model (Bradley & Terry, 1952), is given by:
$\mathbb{P}(y_{w}>y_{l}\mid x)=\frac{\exp(r_{\phi}(x,y_{w}))}{\exp(r_{\phi}(x,y_{w}))+\exp(r_{\phi}(x,y_{l}))}$
To optimize $r_{\phi}$, we minimize the negative log-likelihood of this model:
$L_{R}(r_{\phi})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$
Once $r_{\phi}$ is fine-tuned, it substitutes the initial reward function $r$ and is integrated directly into the reinforcement learning framework, enhancing the model’s performance through explicit optimization via Proximal Policy Optimization (PPO) (Schulman et al., 2017):
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim p_{\nu}}\left[r_{\phi}(x,y)-\beta D_{KL}(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x))\right]$
Here, $\beta$ adjusts the deviation from the base reference policy $\pi_{\text{ref}}$, ensuring the model adheres closely to desired behaviors.
Offline preference optimization. We experimented with Direct Preference Optimization (DPO), which aligns language models with human preferences without the need for an explicit reward model. DPO reparameterizes the reward function $r$ using the following expression:
$r(\mathbf{x},\mathbf{y})=\beta\log\frac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})}+\beta\log Z(\mathbf{x})$ (6)
where $Z(\mathbf{x})$ is the partition function. The objective for DPO then becomes:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(\mathbf{y}_{w}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}_{w}|\mathbf{x})}-\beta\log\frac{\pi_{\theta}(\mathbf{y}_{l}|\mathbf{x})}{\pi_{\text{ref}}(\mathbf{y}_{l}|\mathbf{x})}\right)\right],$ (7)
where $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ are preference pairs consisting of the prompt, the winning response, and the losing response from the preference dataset $\mathcal{D}$. This formulation allows DPO to optimize directly based on preferences without a reward model. We apply LoRA fine-tuning (Hu et al., 2021) for both algorithms to efficiently update model parameters.
Appendix F Human Study Details
F.1 Additional Results
Figure 13: The policy trained using the proposed RLHS outperforms that of RLHF in both true utility (left) and hindsight rating (right). In both plots, each point represents the mean ratio for a scenario, with lines indicating the standard deviation. The identity line is plotted in dashed grey.
F.2 User Interface
In this subsection, we display the interface used in our human study.
Figure 14: Example of user interaction interface for our main human experiments studying the misalignment of RLHF and the effecitveness of RLHS.
Figure 15: Example of user interaction interface for additional human experiments assessing the alignment of LLM actions and feedback with those of humans.
F.3 Participants and data collection
The human subjects were chosen from a high quality Prolific participant pool, where participants were pre-screened to have an approval rate of 95-100 over at least 100 previous submissions. Participants were located in the USA. To assign subjects to experimental conditions, we used random assignment, and each participant was only assigned to one shopping scenario (either one purchasing decision or comparing between two AI shopping assistants). As a negative experience could bias participants’ perceptions of AI chatbots, we ensured that they were not able to retake the study.
The expected duration of the study was 5 minutes, and actually completed the study at a median time of 4:54. Subjects were compensated \$1.10 for their participation, resulting in a hourly wage of \$13.47/hour, which was substantially higher than minimum wage. In addition to participant satisfaction ratings or preferences, participants were asked to provide a brief 2-sentence explanation to explain their ratings or preferences. We manually reviewed these explanations for all participants, and participants that did not provide a reasonable 2-sentence explanation had their data removed from the study. We also removed participants that finished the study in an unreasonably short time ($<$1:30 out of the estimated 5 minutes). Other than this, no data was removed.
This study received IRB approval at [redacted] institution with the record number [redacted].
F.4 Additional human study on alignment between human and AI feedback
We conducted an additional human study to assess how closely the feedback and actions of our AI proxy (Llama-3.1-70B) align with those of human participants. In the study, participants interacted with chatbots from two different stores, taking actions such as purchasing items or leaving the store based on the conversations. After engaging with both stores, participants were asked to choose which store they preferred. We randomly selected 10 scenarios from our training set, with 30 different participants evaluating each scenario. To determine the human preference for each scenario, we employed majority voting. This method was used to ensure that the aggregated choice reflected the consensus among participants, minimizing the impact of individual variability and providing a more robust measure of overall preference. Our analysis revealed that the matching accuracy between LLM-generated feedback and human feedback reached $100\%$. Furthermore, the actions taken by the LLM matched those of human participants with $95\%$ accuracy. These findings suggest that our simulated feedback and actions align strongly with real human behavior.
Appendix G Prompts
In this section, we provide example prompts that we used to generate feedback from AI.
Table 7: Prompt for asking the customer’s satisfaction.
How satisfied are you with the service? Please provide a rating from 1 to 5, based on the following criteria:
1: Very Dissatisfied
   - The chatbot provided misleading or incorrect information.
   - The chatbot’s responses were irrelevant or off-topic.
   - The interaction was frustrating, and my questions were not addressed.
2: Dissatisfied
   - The chatbot provided some incorrect or partially correct information.
   - The chatbot’s responses were occasionally irrelevant or not helpful.
   - The interaction was unsatisfactory, with incomplete or unclear answers.
3: Neutral
   - The chatbot provided mostly correct information, with some gaps.
   - The chatbot’s responses were somewhat helpful but not comprehensive.
   - The interaction was average, neither particularly good nor bad.
4: Satisfied
   - The chatbot provided accurate and helpful information.
   - The chatbot’s responses were relevant and mostly thorough.
   - The interaction was smooth, addressing most of my questions effectively.
5: Very Satisfied
   - The chatbot provided clear, accurate, and highly helpful information.
   - The chatbot’s responses were comprehensive and directly addressed my needs.
   - The interaction was excellent, resolving my queries effectively and efficiently.
Please rate your experience from 1 to 5 based on the above criteria and explain the reasons for your rating.
Table 8: Prompt for asking the customer to make decision.
Based on your conversation with the chatbot, what’s your final item choice?
A) Buy Option A
B) Buy Option B
C) Buy Option C
D) Do not buy anything
Please answer with a capital letter (A, B, C, or D) and explain why.
Table 9: Prompt for generating feedback from AI.
You are a customer seeking to make a purchase, and you visit two different online stores, both of which have a chatbot that you can consult to help you choose.
Both stores contain three options on {item}:
Option A: {price_A}
Option B: {price_B}
Option C: {price_C}
Your absolute requirement is a {item} that must include the following features: {requirement}.
Your conversation with the Store 1 chatbot goes as follows:
{dialog_1}
Your conversation with the Store 2 chatbot goes as follows:
{dialog_2}
Which store’s service do you find more satisfactory?
Appendix H Discussion
H.1 Additional Related Work
Recent work (Lang et al., 2024) shows how partial observability can incentivize deception in RLHF. This is distinct from the problem of human misprediction we address. In their setting, user utility is confined to the immediate time frame of the interaction and does not consider the long-term repercussions on the user’s behavior or well-being after the interaction concludes. Their analysis primarily highlights scenarios where an AI system is incentivized to withhold information to avoid negative feedback scores but does not delve into the real-world impact such deception has on user utility. In contrast, our approach specifically examines the human user’s decision-making process after interacting with the AI system, emphasizing how misalignment or deceptive behavior directly affects their realized utility. We argue that careful consideration of the downstream consequences of human-AI interactions is essential for achieving genuine human-AI alignment.
H.2 Additional limitations and future works
Personalized Hindsight Simulation. Users inherently differ in preferences, risk tolerances, and expertise, causing identical outcomes to have varied perceived utilities. Integrating personalized user models into RLHS could significantly enhance alignment by tailoring simulated hindsight outcomes more closely to individual user objectives. Future studies could explore personalization techniques, leveraging explicit preference elicitation or implicit user behavior modeling to further improve the utility and acceptability of RLHS-aligned systems.
H.3 Broader Impact.
Human evaluators in RLHF often lack full knowledge of AI systems’ internal processes and can misjudge downstream outcomes. This issue makes robust alignment practically challenging to achieve with both closed-source (e.g., ChatGPT) and open-source models, as evidenced by the ever-growing body of literature on FM hallucination, sycophancy, and jailbreak vulnerability. We expect that the introduction of hindsight simulation as a general mechanism for feedback elicitation will make a positive impact by helping inhibit the emergence Goodhart’s law dynamics. We expect the hindsight simulation mechanism to scale favorably as the capabilities of generative AI systems continue to advance in the coming years: the more accurate and powerful the predictive world models leveraged by the AI system in sampling plausible futures when eliciting evaluator feedback, the better-grounded this feedback can be expected to be. This is crucial because increases in capability do not generally grant improvements in alignment; in contrast, RLHS directly takes advantage of highly capable (not necessarily aligned) AI world models to improve the reliability and scalability of value alignment.
Algorithm 1 Human Feedback Loop for RLHS
1:  Step 0: Initialization
2:  ${s}_{0},{z}^{\textit{H}}_{0},{\theta}^{\textit{H}},{o}_{0}^{H}\leftarrow\text{sample\_initial\_conditions}(\mathcal{S},\mathcal{Z}^{\textit{H}},\Theta^{\textit{H}})$
3:
4:  Step 1: AI Prompt Sampling
5:  ${s}_{0}^{\textit{AI}},{o}_{0}^{\textit{AI}}\leftarrow\text{sample\_AI\_prompt}(\mathcal{Z}^{\textit{AI}},\mathcal{O}^{\textit{AI}})$
6:
7:  Step 2: AI Policy Evaluation
8:  Query the AI policy for an action: ${o}^{H}_{1}:=a^{\textit{AI}}_{0}\sim{\pi}^{\textit{AI}}(\cdot\mid{s}_{0},{z}^{\textit{H}}_{0})$
9:
10:  Step 3: Hindsight
11:  for $t=1$ to $T+N$ do
12:     Sample action: $a_{t}\leftarrow\text{sample\_action}({\pi}^{\textit{AI}})$
13:     ${s}_{t+1},{o}_{t+1}^{H}\leftarrow f({s}_{t},a_{t},{o}_{t}^{\textit{H}})$
14:  end for
15:
16:  Step 4: Query Feedback
17:  Query human feedback on the AI policy: $\hat{{U}}^{H}(\pi^{AI})\leftarrow\text{query\_human\_feedback}(\pi^{AI})$
18:
19:  Output or Process Feedback
[…]
H.4 Computing Resources
All experiments were conducted using Nvidia L40 GPUs (48GB memory). A single GPU suffices for inference and LoRA fine-tuning of Llama-3-8B and Llama-2-7B. However, inference with Llama-3.1-70B (used as the judge) requires four GPUs. Fine-tuning typically takes 1–2 days, inference completes within one day, and generating the complete preference dataset for fine-tuning requires more than two days.
Appendix I Additional Qualitative Results
In this section, we provide additional results comparing the qualitative differences between the outputs of policies trained with RLHF and RLHS. We also show a failure case here.
Figure 16: Qualitative results for Llama-2-7b trained with DPO using immediate feedback versus partial hindsight. The model trained with immediate feedback falsely claims that Option B is most affordable with 8K resolution, which is incorrect. In contrast, the model trained with partial hindsight truthfully states that option C is the most affordable option that includes 8K resolution.
Figure 17: Qualitative results for Llama-3-8b trained with DPO using immediate feedback versus partial hindsight. The model trained with immediate feedback falsely claims that Option C can play 3D movies, which is incorrect. In contrast, the model trained with partial hindsight accurately states that Option C’s 3D capability is not specified, and recommends Option B, the cheapest option that includes 3D capability.
Figure 18: Failure case for Llama-2-7b trained with DPO using partial hindsight. The model trained with immediate feedback deceives about specific features, while the model trained with partial hindsight withholds some information. This reveals shortcomings of partial hindsight, as it does not have observations for all other items. Consequently, it might still encourage the agent to deceive about the price or conceal price information.